### v1

# Large Language Models and Causal Inference in Collaboration: A Comprehensive Survey

Abstract
Causal inference has shown potential in enhancing the predictive accuracy, fairness, robustness, and explainability of Natural Language Processing (NLP) models by capturing causal relationships among variables. The emergence of generative Large Language Models (LLMs) has significantly impacted various NLP domains, particularly through their advanced reasoning capabilities. This survey focuses on evaluating and improving LLMs from a causal view in the following areas: understanding and improving the LLMs’ reasoning capacity, addressing fairness and safety issues in LLMs, complementing LLMs with explanations, and handling multimodality. Meanwhile, LLMs’ strong reasoning capacities can in turn contribute to the field of causal inference by aiding causal relationship discovery and causal effect estimations. This review explores the interplay between causal inference frameworks and LLMs from both perspectives, emphasizing their collective potential to further the development of more advanced and equitable artificial intelligence systems.
1 Introduction
Recently Large Language Models (LLMs) have showcased remarkable versatility across a spectrum of critical tasks. An LLM is adept at tasks such as copywriting, enhancing original sentences with their distinct style and voice, responding to knowledge base queries, generating code, solving mathematical problems, and performing classification or generation tasks tailored to user requirements. Moreover, there has been a recent expansion into multi-modal variants, such as Large Visual Language Models (LVLMs) or Large Multi-modal Language Models, which broaden their input/output capabilities to encompass various modalities. This evolution has significantly enhanced both the potential and range of applications of these models.
In this survey, our primary focus is on Transformer-based Large Language Models (LLMs). The capability of LLMs is fundamentally rooted in their inference abilities, which dictates their proficiency in comprehending, processing, and providing solutions to various inquiries, as well as their adaptability to societally impactful domains. Consequently, extensive research efforts have been dedicated to measuring and enhancing these capabilities, ranging from assessing the reasoning abilities of LLMs to scrutinizing their decision-making processes and addressing challenges such as concept alignment across different modalities and mitigating hallucination. In addition, since LLMs are trained on extensive human knowledge with billions of parameters, they sometimes face challenges in appropriately prioritizing or downplaying what they have learned in different scenarios. This can lead to issues such as domain shift, where the model’s performance degrades on data that differ from the training set, and long-tail bias, where infrequent examples are not handled as effectively.
In many instances, language tasks require not only predicting or generating text based on patterns in the data but also understanding the underlying causal mechanisms driving these patterns. Causal inference has shown great potential in improving predictive accuracy, fairness, robustness, and explainability of Natural Language Processing (NLP) models [23]. With the advent of generative LLMs, a significant transformation has occurred across various NLP fields, attracting increased research interest in applying causal inference to address LLM-related challenges and augment their functionality. Such transformation also motivates this survey to outline causal methodologies and their implementation in LLMs, emphasizing their role in enriching our comprehension and application of language models.
{forest}
for tree= draw=red, align=center, anchor=west, child anchor=west, grow’=0, rounded corners, font= [Causal and Large Language Models (LLMs), rotate=90, anchor=south, text width=5.5cm [Causal for LLMs, [Reasoning Capacity [Model Understanding [[109, 79, 43, 47, 1, 88, 116], fill=yellow!20] ] [Commonsense Reasoning [[27, 98, 114, 123, 62, 17], fill=yellow!20] ] [Counterfactual Reasoning [[11, 70, 50, 60], fill=yellow!20] ] ] [Fairness and Bias [[68, 2, 82, 22, 124, 95, 65, 37], fill=yellow!20] ] [Safety [[15, 9, 122, 118, 69, 56], fill=yellow!20] ] [Explanation [[35, 10, 33, 119, 55], fill=yellow!20] ] [Multi-modality [[74, 45, 49, 85, 86, 121], fill=yellow!20] ] [Benchmark [[103, 115, 36, 107, 41, 72, 40, 85, 49, 18, 27], fill=yellow!20] ] ] [LLMs for Causal [Treatment Effect Estimation [[19, 24, 113], fill=yellow!20] ] [Causal Relationships Discovery [[44, 71, 4, 5, 61, 92, 7, 91, 8, 7], fill=yellow!20] ] ] ]
Figure 1: Structure Overview
Meanwhile, this survey also aims to explore how LLMs can help with the causal inference framework. Causal inference is formally defined as an intellectual discipline that considers the assumptions, study designs, and estimation strategies that allow researchers to draw causal conclusions based on data [75]. Causal inference has three main origins: potential outcomes, graphs, and structural equations, each serving unique purposes [110]. The potential outcomes framework [75] focuses on estimating causal effects through statistical inference and treatment comparisons. Graphical models, meanwhile, excel in mapping out causal pathways and visualizing relationships, with nodes representing variables and edges denoting directional influences. In this survey, we mainly discuss Pearl’s formulation of causal graphs [76], which formalized causal graphical models for presenting conditional independence relations among random variables using directed acyclic graphs (DAGs).
We summarize how LLMs can help causal inference in its two important components, i.e., causal relationship discovery and treatment effect estimation. Determining causal relationships among variables is a fundamental step in a causal inference framework because the estimation of the causal effect of variable $A$ on variable $B$ requires causal assumptions on causal relationships of other variables with $A$ and $B$. Traditionally, researchers rely on experts with subject matter knowledge to set the ground for such causal relationships. Causal discovery methods [108] provide an alternative by discovering causal graphs from observational data. LLMs have demonstrated abilities to determine such causal relationships based on pre-trained knowledge or given text. They can also be integrated with causal discovery methods to further increase the reliability of the outcome. Estimating treatment effects is central to causal inference but is hindered by the absence of counterfactual data in many cases. Leveraging the strong counterfactual reasoning abilities of LLMs, researchers have developed various ways to generate high-quality counterfactuals to enable treatment effect estimation.
The structure of the survey is given in Figure 1. We start with an introduction to recent advancements in Large Language Models in Section 2. Then we provide an overview of causal inference methods that are used and can be used to improve LLMs in Section 3. In the first half of the paper we go through how these methods are used in various problems in the LLM community: Section 4.1 overviews how causal methods are used to measure and improve the reasoning capacity of LLM, Section 4.2 and Section 4.3 focus on the fairness and safety issues while Section 4.4 introduces how the explainability of LLM are approached by causal inference methods. We also discuss the extension to construction and development of multi-modality large models in Section 4.5. Finally, we list existing work on evaluation and benchmarking LLMs from a causal perspective in Section 4.6. In the second half of the survey, we shift to how LLM extends the boundary of causal inference. Section 5.1 explains current assumptions, limitations and bottlenecks of causal inference. Section 5.3 and Section 5.2 states current works on improving treatment effect estimation and causal discoveries. We highlight several future directions in Section 6.
2 Background of Large Language Models
Large Language Models (LLMs) have transformed the way we interact with and process language, opening up new possibilities for natural language understanding, generation, and communication [34]. These models are in a constant state of evolution, continually expanding the limits of what is achievable in language processing and artificial intelligence [73]. In this paper, we mainly focused on Transformer-based LLMs, and we provide an overview of their recent progress in this session.
The major breakthrough of Large Language Models (LLM) came in 2017 when the Transformer [93] was introduced. In a groundbreaking shift, the Transformer technology emerged, mastering the art of grasping long-term linguistic connections. This innovation wasn’t just theoretical—it allowed for simultaneous training on multiple GPUs, paving the way for the creation of significantly larger models [16]. Then came a pivotal moment in 2018: the unveiling of OpenAI’s GPT-1 [78]. This was more than just another step in natural language processing (NLP); it was a leap, thanks to its transformer-based architecture. Boasting 117 million parameters, GPT-1 wasn’t merely crunching data; it was weaving contextually coherent sentences, showcasing the transformative power of transformers in redefining NLP tasks [3]. Despite its initial limitations, GPT-1 didn’t just make a mark; it laid the groundwork for a new wave of AI exploration and sparked intense competition in the realm of LLMs.
In 2020, OpenAI released GPT-3 [20], which was able to generate highly coherent and natural-sounding text [13]. It was a big deal because it showed just how awesome these huge language models could be at a bunch of different language tasks [26]. Riding on the high from GPT-3’s success, OpenAI released the next iteration of their language model, GPT-4 [2, 67, 52] with the ability to generate even more coherent and natural-sounding text. Following GPT-4’s success, Google wasn’t far behind with their Bard [66]. Amazon jazzed up Alexa [42] with some cool AI features, Huawei jumped in with their Pangu models [111] and Alibaba proposed their QWEN models [6]. Then Meta brought out this thing called LLaMA [90], which was all about the first open-source foundation models. Compared to LLaMA, LLaMA2 [90] has made more explorations in reinforcement learning from human feedback (RLHF) and developed a chat-oriented version called LLaMA-chat, which generally outperforms existing open-source models across a range of helpfulness and safety benchmarks. A large number of researchers have extended LLaMA models by either instruction tuning [57] or continual pretraining. Alpaca [89] is the first open instruct-following model fine-tuned based on LLaMA. In addition, Vicuna [77] is another popular LLaMA variant, trained upon user-shared conversations collected from ShareGPT [31].
As diverse versions of LLMs emerge, they encounter common challenges. In this survey, we demonstrate that many of these challenges can be effectively addressed through causal methods. These methods encompass enhancing the reasoning capacity of LLMs, tackling fairness concerns and mitigating potential biases, ensuring safety, and enhancing the explainability of model outputs, and their extension to multi-modality versions.
Building upon this progress, there is now a burgeoning interest in expanding the scope of these models to encompass visual data, giving rise to the emergence of Large Visual Language Models (LVLMs). These models aim to integrate the understanding of both textual and visual information, opening up new avenues for more comprehensive AI systems capable of interpreting and generating content in multimodal formats.
One of the most prevalent approaches is inserting visual features as supplementary inputs to LLMs and aligning them with textual features. This method has been adapted in several large vision-language models (LVLMs) such as MiniGPT-4 [127], LLaVA [59], Mplug-Owl [106] and so on [55, 57, 117, 96, 106, 58, 51]. In this survey, we also demonstrate that the causal methods could help concur challenges encountered in existing LVLMs as above.
3 Brief Introduction of Causal Inference
In this section, we present the background knowledge of causal inference, including task descriptions, basic concepts and notations, and general solutions.
Generally speaking, the task of causal inference is to estimate the causal relationship among variables. The variables of interest are referred to as treatment, naturally, the effects of treatments are referred to as treatment effects. For example, suppose two treatments can be applied to patients: Treatment Plan A and B. When A is applied to a certain patient cohort, the recovery rate is $70\%$ while when B is applied to the same cohort, the recovery rate is $80\%$. The change of recovery rate is the effect of that treatment assets on the recovery rate.
Ideally, the treatment effect can be measured as follows: applying different treatments to the same cohort, and then the difference in the effect is the treatment effect. However, in real-world scenarios, this ideal situation because it is impracticable for perfectly controlled experiments in most cases. For example, in the above case, you can only apply one treatment to the same cohort at the same time. In reality, an alternative is to conduct random controlled trials, in which the treatment assignment is controlled, such as a completely random assignment. In this way, the groups receiving different treatments can be used to measure the difference in effect. Unfortunately, even performing randomized experiments is expensive, time-consuming, and may cause ethical concerns in some cases. Therefore, estimating the treatment effect from observational data has attracted growing attention due to the wide availability of observational data, and methods are developed for the investigation of the causal effect of a certain treatment without performing randomized experiments.
3.1 Potential Outcome Framework
One of the most influential frameworks in identifying and quantifying causal effects in observational data is the potential outcomes framework [80]. The potential outcomes approach associates causality with manipulation applied to units, and compares causal effects of different treatments via their corresponding potential outcomes. Following [80], we state basic concepts in the potential outcome framework.
Unit. A unit is the atomic research object in the treatment effect study. A unit can be a physical object, a firm, a patient, a person, or a collection of objects or persons, such as a classroom or a market, at a particular time point [80]. Under the potential outcome framework, the atomic research objects at different time points are different units.
Treatment. Treatment refers to the action that applies (exposes, or subjects) to a unit. For each unit-treatment pair, the outcome of that treatment when applied to that unit is the potential outcome.With N treatments $T=\{1,2,3,...,N\}$, the potential outcome of applying treatment $T_{i}$ is denoted as $Y(T=T_{i})$. The observed outcome is the outcome of the treatment that is actually applied. And the counterfactual outcome is the outcome if the unit had taken another treatment.
Treatment Effect The treatment effect can be quantitatively defined using the above definitions. The treatment effect can be measured at the population, treated group, subgroup, and individual levels. At the population level, the treatment effect is estimated as the Average Treatment Effect (ATE). At the subgroup level, the treatment effect is called the Conditional Average Treatment Effect (CATE).
Definition 3.1 (Binary Average Treatment Effect(ATE)).
Suppose we want to measure the treatment effect of a treatment $T=1$. Then the average treatment effect is defined as:
$\mathbb{E}[Y(T=1)-Y(T=0)]$ (1)
where $Y(T=1)$ and $Y(T=0)$ denote the potential treated and control outcome of the whole population respectively.
Definition 3.2 (Conditional Average Treatment Effect (CATE)).
$\mathbb{E}[Y(T=1)|X=x]-\mathbb{E}[Y(T=0)|X=x]$ (2)
where $\mathbb{E}[Y(T=1)|X=x]$, $\mathbb{E}[Y(T=0)|X=x]$ are the potential treated and control outcome of the subgroup with $X=x$.
At the individual level, the treatment effect is defined as Individual Treatment Effect (ITE). In some literature, ITE is treated as the same as CATE [75].
3.2 Causal Graphical Models
The potential outcome framework is powerful in recovering the effect of causes. In a potential outcome framework, causal effects are answered by specific manipulation of treatments. However, when it comes to identifying the causal pathway or visualizing causal networks, the potential outcome model has its limitations. In the front of the challenge, causal graphical models utilize directed edges to represent causalities and encode conditional independence among variables in the graphs.
3.2.1 Structural Equation Models (SEMs)
One of the most widely-spread formulations is the Structural Equation Model [99, 76], where linear structural equation models are used to present causal relationships by directed edges, which differentiate correlation from causation when the graph structure is given. The linearity assumption was later been relaxed by [76] and it formalized causal graphical models for presenting causal relations using Directed Acyclic Graphs (DAGs).
Specifically, consider the random variable $\mathbf{X}\in\mathcal{R}^{D\times N}=[X_{1},X_{2},...,X_{N}]$, the linear SEM consists of a set of equations of the form:
$X_{i}=\beta_{0i}+\sum_{j\in pa(X_{i})}\beta_{ji}X_{j}+\epsilon_{i},\quad i=1,2,3,...,N$ (3)
where $pa(X_{i})$ denotes the set of variables that are direct parents of $X_{i}$. $\epsilon_{1},\epsilon_{2},...,\epsilon_{N}$ are mutually independent noise terms with zero mean, $\beta_{ji}$ are coefficients that quantify the causal effect of $X_{j}$ on $X_{i}$.
While the non-parametric SEM takes the form:
$X_{i}=f_{i}(\mathbf{X}_{pa(i)},\epsilon_{i}),\quad i=1,2,3,...,N$ (4)
The random variables $\mathbf{X}$ that satisfies the model structure of the form in Equation 3 or Equation 4 can be represented by a directed acyclic graph (DAG) $G=(V,E)$, where $V$ is the set of associated vertices, each corresponding to one of a variable of interest $X_{i}$, and $E$ is the corresponding edge set.
With pre-specified DAG and assumptions on the latent variables, the coefficients between the latent variables are identifiable[46].
3.2.2 Bayesian Network
Causal inference can be naturally embedded in graphical model frameworks since the dependencies and interactions between variables can be presented by graphs with probabilistic distributions, in which nodes correspond to variables of interest and edges represent associations. One general solution except for SEMs is to use a Bayesian Network to represent the causal relationship.
In Bayesian networks, causalities among variables are represented in the form of DAGs with directed edges carrying causal information.
A joint probability distribution $\mathbb{P}$ factorizes with respect to a DAG $\mathcal{G}$ if it satisfies:
$f(X_{1},X_{2},...,X_{N})=\prod_{i}f(X_{i}|\mathbf{X}_{pa(i)})$ (5)
In the next section, we show a comprehensive survey of how existing works help with the tasks and challenges in LLMs in detail.
4 Causal Inference for Large Language Models
LLMs can significantly benefit from causal inference as it enhances their ability to understand and reason about cause-and-effect relationships within data. In this section, we review how LLMs can benefit from a causal view in the following perspectives, understanding and improving the LLMs’ reasoning capacity (Section 4.1), addressing fairness (Section 4.2) and safety (Section 4.3) issues in LLMs, complementing LLMs with explanations (Section 4.4), and handling multimodality (Section 4.5). We then organize benchmark datasets from these perspectives in Section 4.6.
4.1 Reasoning Capacity
4.1.1 Model Understanding
LLMs have demonstrated many emerging abilities in language generation and certain reasoning tasks [14, 44]. As the reasoning process is often associated with causal factors, it is logical to first understand and evaluate LLMs’ reasoning ability from a causal lens. Zečević et al. [109] argued LLMs are not causal and hypothesized that LLMs are simply trained on the data, in which causal knowledge is embedded. Thus, in the inference stage, the LLMs can directly recite the causal knowledge without understanding the true causality in the context. Similar behaviors are exhibited in a Causal Reasoning Assessment Benchmark, CRAB, that consists of 1.8K causal frames and 352 causal chains in real-world narratives [79], where LLMs are required to output the causality class (high, medium, low, or no causality) between variables. They show that LLMs can capture explicit causal statements in pre-training data, but face performance drop when applying causal reasoning to new distributions, i.e., events that happened after the pre-training phase. Kim et al. [43] examined LLMs’ abilities in understanding the causalities of both scientific papers and newspapers. Their evaluation protocol is designed for the LLM to tell whether a statement is causal, conditional causal, correlational, or no relationship. The results show that ChatGPT performs worse than a fine-tuned BERT model in understanding causality. Abdali et al. [1] show the effectiveness of applying LLMs to diagnose the cause of issues from Microsoft Windows Feedback Hub. Li et al. [47] showed that LLMs can identify dynamical (spatio-temporal) effects. However, how to infer the relationship and interactions of them is still challenging for LLMs, which are more emphasized as causal structures in causal inference.
Another important line of work is to understand LLMs’ hallucination and faithfulness in knowledge reasoning by considering causal effects. Tang et al. [88] proposed a multi-agent system, CaCo-CoT, where some LLMs are reasoners and others are evaluators. Reasoners try to provide causal solutions, while evaluators try to challenge the reasoners with counterfactual candidates. With the cooperative reasoning framework, CaCo-CoT helps to improve causal-consistency. Zhang et al. [116] identified the potential knowledge bias pretrained in the LLMs as the confounder which causes incorrect answers and hallucinations. Zhang et al. [116] proposed a chain-of-question framework to generate sub-questions necessary to answer a question, and involve humans in the loop to provide the correct answers. With human annotation in the loop, the confounding causal effect can be reduced, thus mitigating the spurious correlation.
4.1.2 Commonsense Reasoning
Commonsense reasoning involves the ability to apply everyday knowledge and intuitive understandings of the world to make decisions or draw conclusions, which is vital for LLMs’ contextual understanding and human-like interactions [21, 84]. This section briefly summarizes the commonsense reasoning ability of LLMs under various settings [27, 98] and the employment of causally motivated methods in improving commonsense causality reasoning [114, 123, 62, 17, 120].
Through the evaluation of ChatGPT’s performance on event causality identification, causal discovery, and causal explanation generation, Gao et al. [27] have shown that ChatGPT is not a good reasoner but a good causal explainer. ChatGPT, and even gpt-4 is outperformed by baseline methods based on fine-tuned small pre-trained language models on event causality identification and does poorly on causal discovery. They also observe serious hallucinations on causal reasoning under In-Context learning and Chain-of-Thought settings A similar conclusion has also been obtained in [98], which analyzed the causal question answering capabilities of LLMs. It concluded that LLMs did not arrive at their answers through reasoning but through memorization of the corresponding question and answer pair. Even if ChatGPT may not have causal reasoning ability, it generates accurate and detailed causal explanations in some cases [27].
To improve the commonsense causality reasoning of LLMs that identify causes from effects in natural language, ROCK (Reason O(A)bout Commonsense K(C)causality) [114] balances confounding effects using temporal propensities. The central question in this framework is the estimation of Average treatment effect (ATE). Given two events such that $E_{1}$ precedes $E_{2}$, the strength of causation from $E_{1}$ to $E_{2}$ can be estimated by the change in the likelihood of $E_{2}$’s occurrence due to intervening on $E_{1}$, denoted as ATE. While ROCK adopts a potential-outcome framework, Chen et al. [17] uses a conversation cognitive model based on intuition theories, and transforms intuitive reasoning into a structural causal model. Aiming to improve conversation reasoning, the authors incorporate mental states such as desires, memory, experience and emotion, as an unobservable exogenous variable that implicitly affects the corresponding observable utterance:
$H=AH+E,$ (6)
where $H$ is the embedding of the utterance, $E$ is the exogenous variable, and $A$ is the adjacency matrix.
Other than facilitating reasoning ability of LLMs directly, Zheng et al. [123] use causal inference to preserve commonsense knowledge from Pre-trained language models for fine-tuning to prevent catastrophic forgetting; Lu et al. [62] focus on improving LLM’s ability in generalized procedural planning with commonsense-infused prompts. In particular, Zheng et al. [123] abstract the fine-tuning process as a causal graph and discovered that catastrophic forgetting is due to missing causal effects from the pre-trained data. To preserve old knowledge from a pre-trained language model, an objective for fine-tuning (causal effect tuning) is introduced. For procedural planning tasks, Lu et al. [62] proposed to learn causeeffect relations among complex goals and stepwise tasks, and reduced spurious correlation among them via front door adjustment.
4.1.3 Counterfactual Reasoning
Another potential use of LLMs’ generative capacity is to generate counterfactuals for data-augmentations for small language models. Given a text $x$ and a black-box classifier $B$, the counterfactual text $\tilde{x}$ of text $x$ should satisfy the following requirement [11, 70]:
  1. 1.
    $\tilde{x}$ has a different class than $x$, $B(x)\neq B(\tilde{x})$.
  2. 2.
    $x$ and $\tilde{x}$ differ only by minimal lexical changes.
  3. 3.
    $\tilde{x}$ is a feasible text and the commonsense constraint is satisfied.
This section discusses briefly the performance of LLMs in generating counterfactuals [50, 60] and the effort of improving their qualities [70].
Li et al. in [50] examined the effectiveness of LLMs’ performance in generating counterfactuals under four tasks: (1) sentiment analysis (SA) that alters sentiment polarity; (2) natural language inference (NLI) that given a premise sentence and a hypothesis sentence, make a change to the hypothesis sentence to alter the relationship between it and the premise sentence; (3) named entity recognition (NER) that changes the entities in a sentence whose type is the same as the original entity type; (4) relation extraction (RE) that changes the relation between the head and tail entity. It has been demonstrated that for simple tasks like SA and NLI, data augmented via LLMs can mitigate potential spurious associations. For more complicated tasks like RE, LLMs may generate low-quality counterfactuals. In order to generate high-quality counterfactual, detailed instructions are crucial. However, counter-intuitively, chain-of-thought doesn’t always help as it may even lead to significant performance decreases under some settings. On the other hand, Liu et al. in [60] evaluated abductive reasoning and counterfactual reasoning abilities for large language models of code (Code-LLMs) and compared them with text models. With code-prompts designed to tackle causal reasoning tasks, it has been shown in [60] that Code-LLMs achieve better results than text models.
As shown in the evaluation [50], though LLMs as counterfactual generators can enhance the performance of small language models on simple tasks such as sentiment analysis and natural language inference, the generated counterfactuals fail to have any significant effect on complex tasks like relation extraction. Miao et al. in [70] claim that this is due to the failure of identifying causal terms correctly and ignoring the commonsense constraint. To amend this, they proposed a framework to generate commonsense counterfactuals for stable relation extraction via an intervention-based strategy. This framework is demonstrated to have enhanced the stability on relation extraction tasks under various settings including the low-resource, out-of-domain and adversarial-attack scenarios.
4.2 Fairness and Bias
Fairness and bias are pivotal factors in deploying language models effectively and ethically. Bias is common in pretrained language models as they capture and potentially amplify undesired social stereotypes and biases [68, 2, 126, 100, 96]. An example of bias in language models includes gender associations with specific professions, such as male firefighters and female nurses [82]. Causality-based methodologies offer a promising approach for mitigating biases in language models by discerning the origins of bias through a causal perspective. Bias mitigation is then followed by eliminating the unwanted spurious correlation between generative factors though different types of causal intervention or causal invariant learning.
Ding et al. [22] introduce a proxy variable related to gender bias in the causal graph, and use two different ways to eliminate the potential proxy bias and unresolved bias under the linear structural equation model. Zhou et al. [124] believe that the backdoor path between the ground truth label and the non-causal factors is the source of bias, and uses the Independent Causal Mechanism (ICM) principle to mitigate bias. Their proposed method, Causal-Debias, achieves causal intervention and creates interventional distributions with respect to different demographic groups by augmenting and expanding the original data distribution. Wang et al. [95] raise concerns over the precision of parameter estimation in existing causal models and introduce several intermediate variables that are the causal children of raw text and the causal parent of the input to language models. Under this assumption, they propose to eliminate bias by performing a ‘do’ operation on the intermediate variables for both white-box LLMs and black-box LLMs. Madhavan et al. [65] consider the tokens generated by generative language models trained with causal language modeling (CLM) objectives as a causal graph, and analyze the bias under this model. Jenny et al. [37] use Activity Dependency Networks (ADNs) to describe the causality effect between normative variables, such as clarity and authenticity, to structure the cause of bias. The authors argue that using ADNs can better explain previously simplified views of bias using just correlation, and display the complexity nature in identifying and mitigating biases in large language models.
4.3 Safety
With the application of language models in multiple downstream tasks, researchers have observed the unreliability phenomenon of LLMs in knowledge probing [15] or downstream inference tasks [9, 122, 118, 69]. There is increasing interest in applying the causal inference technique to analyze the causality of the non-robustness of the model and adjust the treatment to resolve the challenges [15, 118, 122, 56].
Knowledge probing
To elicit the knowledge encoded in LLMs, previous work uses prompt-based prompting, that is, querying LMs with task-specific prompts. However, in this process, LLMs face challenges of unreliability [15, 102], such as using shortcuts to complete the probing and generating different predictions for the semantically equivalent prompt. To explore the reason behind the non-robustness, an empirical study from [15] constructs a structural causal model (SCM) containing 11 variables and identifies three types of bias in the knowledge probing procedure: prompt preference bias, instance verbalization bias, and sample disparity bias. Blocking the corresponding backdoor path for each type of bias effectively eliminates the bias. This finding reveals the potential of constructing an SCM with LLM-related variables to improve LLM robustness.
Downstream Inference Tasks
In addition to knowledge probing, LLMs show vulnerability when encountering attacks on the prompts in downstream inference tasks [118, 122]. By simply translating text tokens in input prompts to emoji sequences, LLMs generate more severe hallucinations [122]. Some neurons within LLMs also have an unreasonably high causal effect on the model response and, by changing the value of that neuron, LLMs will produce meaningless responses [122].
The main reason why LLMs fail in prompt attacks is that it uses the spurious correlation to make an inference [118]. Training LLMs to learn the causal relationship between input $x$ and output $y$ is an intuitive method of better resisting prompt attacks. The randomized smoothing technique [112, 39] can model the interventional distribution $p(y|do(x))$ by assuming discrete adversarial perturbations as the Gaussian distribution [118]. The method of smoothing in the latent semantic space is more robust against known attacks such as word substitutions, paraphrasing, and token position change [118].
Causal inference techniques are useful to induce robustness for LLM applications. For the setting without prompt attacks, mitigating LLM unreliability can be achieved by identifying the causality and blocking the backdoor path. Moreover, smoothing the latent semantic space is effective in resisting prompt attacks. Despite the progress of causality analysis, there are some directions to explore, such as the causal relationship between the pretraining corpus and model robustness and the evaluation method of robustness in the generation task.
4.4 Explainability
Explainability in LLMs refers to the capacity to elucidate how these models arrive at their conclusions, enhancing transparency and trustworthiness in AI decision-making processes [32, 54]. Many work have tried to explain and understand the inner workings of LLMs [35, 10, 33, 119, 55]. We summarize research efforts that probe the causal mechanism in LLMs from the following three directions: intervening the inputs or prompts, intervening inner components of LLMs, and abstracting the working mechanism into a causal graph.
Inputs or Prompt Intervention
Input intervention, as a data-centric method, is to create counterfactual input text by changing the treated feature in the text and to observe the model behavior on the original and counterfactual texts. Recruiting individuals to produce counterfactual texts typically incurs significant expenses. However, the advent of LLMs has shown that generating counterfactual inputs can be achieved with less cost. LLMs can first identify the features in input texts causally associated with the predictions and are capable of changing the identified features to create the counterfactual texts [12, 28]. These counterfactual texts can be utilized to investigate the causality of the LLM and can serve as a training dataset to learn a matching model, where the matched counterfactual pairs have similar embeddings [28].
Various works have developed different prompting methods and found whether the prompting methods are causally associated with the final output of LLMs [81, 125, 97]. However, the causal effect of prompting methods, such as chain-of-thought (CoT), and the final output is ambiguous. Prompt intervention, which alters only one particular aspect of prompts, is proposed to understand the contributions of each component of prompts on model behavior [64]. The experiments from [64, 38] first find that the linguistic features and the grammar have a large effect on the LLM outputs. Then, intervention in intermediate variables in prompts leads to consistent final answers with the expected output of the hypothesized causal model [87]. These findings suggest that LLMs realize the causal model suggested by their CoTs to a high extent, but LLMs also utilize spurious correlations such as sentence length to make responses.
Input or prompt interventions are data-centric methods to probe the LLM mechanism, applied in both open-source or black-box LLMs. However, the detailed information cascade inside the LLMs cannot be discovered by such methods, so intervention on the inner model components is proposed.
Inner Component Intervention
The attention mechanism and multilayer perceptron (MLP) layers are the essential components in the structure of state-of-the-art (SoTA) LLMs. Existing work exchanges the activation values in MLP and attention layers of different inputs to probe the function of MLPs and attention mechanism in generating the answers [83]. The experimental results indicate that LLMs transfer the information of inputs from midsequence early layers to the final token using the attention mechanism. Due to the complexity of LLMs, current work focuses only on math word problems with four fundamental arithmetic operators [83]. It is an interesting direction to generalize the component intervention to other downstream tasks.
Causal Graph Abstraction
An intuitive way to characterize causality within LLMs is to abstract the working mechanism of LLMs into a causal graph. Boundless Distributed Alignment Search (DAS) [101] by replacing brute-force searching the original DAS [29] with learnable parameters, has been effective on the Alpaca model [89]. Given four pre-defined causal models, Boundless DAS extracts two of them as the accurate hypotheses as the abstracted causal graph of the Alpaca model. However, the Boundless DAS method is restricted by the given causal hypothesis, and the future direction can explore how to abstract the causal graph in LLMs without prior causal graphs.
Current explainability work from a causal view use LLMs to interpret the causality in real-world events. LLMs can generate high quality counterfactuals. By altering the inner values of LLMs and abstracting the causality in LLMs, the current work has pointed out a direction to characterize the inner causality of LLMs on various tasks.
However, with these mentioned works in LLM interpretation, the scope of interpretation focuses on tasks that have a clear causal graph between task inputs and outputs such as simple math word problems. Probing the causality in LLMs for more complex generation tasks, such as question answering or summarization tasks, could provide more insight on the inner mechanism of LLMs.
4.5 Multi-modality
Large vision-language models have been popular in many applications [63, 48, 59, 30]. How to conduct causal reasoning on both images and texts can be crucial in correctly answer multimodal questions. Pawlowski et al. [74] examined LLMs’ causal reasoning abilities and showed that the causal knowledge in the language models can be too strong a prior which often causes the model to neglect visual evidence. Ko et al. [45] proposed to alleviate the problem by adding self-consistent generation prediction, in which the three input V, Q and A are individually predicted based on the other two inputs. Specifically, Li et al. [49] proposed a image generation framework with causal reasoning and created a novel VQA datasets whose questions requires causal explanations.
Another important question is to understand the spatial-temporal causal relationship of the visual elements within the images and videos. Su et al. [85] proposed a framework, CaKE-LM, to leverage the pretrained causal knowledge in the language model to understand the causal relations of the events in a video. Based on the generated causal reasoning results, CaKE-LM can further generate the question-answer pairs and construct a new benchmark for causal video question-answering. Tai et al. [86] proposed the link-context learning method to strengthen the LLM’s in-context learning abilities by instructing the model to understand the underlying causal relationship between the demonstration data points. Zhao et al. [121] suggested that there are two types of causal relationships in VQA. They proposed a prompting method, causal context generation, to engage contextual information for better VQA precision.
4.6 Evaluation and Benchmark
In this section, we list existing evaluation metrics and benchmarks from a causal perspective for LLMs as listed in Table 1. The causal evaluation mainly focuses on three aspects: the Model Understanding (MU) ability, the Commonsense Reasoning (CR) ability, the Counterfactual Reasoning (CF) ability, and the Fairness/Debias (FD) ability.
Reference MU CR CF FD Language Multimodal
ECHo [103]
CREPE [115]
CLOMO [36]
IfQA [107]
Cladder [41]
MoCa [72]
CORR2CAUSE [40]
CVidQA [85]
VQAI [49]
Chen et al. [18]
Gao et al. [27]
CRAB [79]
HELM [53]
Fair-Prism [25]
Biasasker [94]
Table 1: We summarize the existing evaluation benchmarks. Based on the evaluation tasks, we categorize the benchmarks into three categories: Model Understanding (MU), Commonsense Reasoning (CR), Counterfactual Reasoning (CF) and Fairness/Debias (FD). Based on the modalities of the data samples, we identify the benchmarks with only textual inputs (Language) and those with multimodal inputs (Multimodal).
The benchmarks in model understanding (MU) focus on evaluating and understanding existing LLMs’ causal reasoning abilities in both natural language [41, 72, 40, 27, 79] and vison-language [85, 49, 103]. In addition, some benchmarks [72, 27] also provide model understanding in comparison with human causal reasoning and moral judgments. Commonsense reasoning benchmarks (CR) evaluate LLMs on tasks that require extensive commonsense knowledge for both textual-only context [115, 41] and multimodal context [103]. Contexts with commonsensical and anti-commonsensical are constructed in [41], to further investigate whether LLMs use averaged-out causal reasoning. Evaluating LLMs’ counterfactual reasoning (CF) abilities is essential in enabling explainable model reasoning and calibration of the generated rationales. Huang et al. [36] introduce a specific task and benchmark for assessing LLMs’ logical counterfactual thinking. Yu et al. [107] contribute a novel dataset to challenge LLMs in counterfactual reasoning in an open-domain QA context. Chen et al. [18] investigate the ability of LLMs to provide explanations that aid in understanding their reasoning process, particularly in the context of counterfactual scenarios. The fairness and bias (FD) evaluations are particularly in addressing biases, fairness, and the overall transparency of language models. HELM [53] is a comprehensive evaluation benchmark including previously neglected areas for fairness. Fair-Prism [25] focuses specifically on fairness-related harms in models, which are identified and measured by detailed human annotations. Biasasker [94] presents an automated framework to identify and measure social biases by probing the models with specially designed questions.
5 Large Language Model for Causal Inference
Causal inference, serving as a potent tool for addressing challenges in LLMs, heavily relies on world knowledge. As previously mentioned, there are three primary origins of causal inference: the potential outcome framework, graph-based causal methods, and the structural equations community. The potential outcome framework relies significantly on several assumptions to facilitate the comparison of treatment effects among groups/individuals. One of the most challenging aspects of applying the potential outcome framework lies in ensuring that these assumptions hold true in reality. In this section, we first examine these assumptions and subsequently illustrate how they are relaxed in existing literature. The graph-based causal methods and structural equation models also necessitate a certain level of understanding of the underlying causal graph. For instance, Directed Acyclic Graphs (DAGs) serve as a fundamental assumption, and many structural equation models assume a degree of linearity or that the input distribution adheres to specific probability distributions. In our review, we also explore how existing methods verify the distribution in the input data and extend current methodologies to accommodate more complex distributions with the assistance of LLMs.
5.1 Fundamental Assumptions in Estimating Treatment Effect
In existing causal inference literature, several assumptions are adopted to estimate the treatment effect. Here we discuss the three most commonly used assumptions and then show how the development of large language models could help relax or challenge these fundamental assumptions.
Assumption 5.1 (Stable Unit Treatment Value Assumption ).
The potential outcomes for any unit do not vary with the treatment assigned to other units, and, for each unit, there are no different forms or versions of each treatment level, which leads to different potential outcomes.
This assumption emphasizes the independence of each unit when estimating the treatment effect that the units do not interact with each other. From a statistical perspective, it is equivalent to assuming each treatment assignment subject is iid.
Assumption 5.2 (Ignorability/Unconfoundedness ).
Given the observable background variable, X, treatment assignment T is independent of the potential outcomes.
The 5.2 states that if the background variable $X$ is the same for two patients, then (1) the treatment assignment should be the same. (2) the potential outcome should also be the same. In other words, these two patients are treated as identical units, thus can be used to estimate the treatment effect if they are assigned different treatments in the static dataset, as the treatment assignment is treated as random.
Assumption 5.3 (Positivity ).
For any value of X, treatment assignment is not deterministic. i.e.,
$P(T=t|X=x)>0,\forall t,x$ (7)
This assumption tries to guarantee that the treatment effect can be estimated and that we can always find a comparable sample. In the binary case, if the task is to estimate the performance of a specific treatment $T=1$, we would need to compare the potential outcome of patients receiving the treatment against those who are not treated, requiring dataset points in both cases.
5.2 Treatment Effect Estimation
One main obstacle of traditional causal methods is the lack of counterfactual data, making the estimation of causal effects a difficult problem in practice. Chen et al. [19] proposed a new method for automatically generating high-quality counterfactual data at scale called DISCO (DIStilled COunterfactual Data). Specifically, it prompts to generate of phrasal perturbations with a large general language model. Then, a task-specific teacher model filters these generations to distill high-quality counterfactual data. In addition, Feder et al. [24] apply treatment effect estimation to align knowledge for generalization towards different domains. Zhang et al. [113] try to optimize the treatment effect estimation on unlabeled datasets by performing self-supervised causal learning through LLMs. Through exploring the primal-dual connection between optimal covariate balancing and self-attention, their method facilitates zero-shot causal inference through the final layer of a trained transformer-type architecture, contributing to a foundation model for treatment effect estimation.
5.3 Causal Relationships Discovery
Discovering causal relationships between variables is a fundamental step in causal inference as it enables the identification and estimation of causal effects. In this section, we introduce papers discussing how LLMs can help discover causal relationships.
One line of this work focuses on casual relationship extraction or causality extraction which extracts causal relationships between two variables from text directly. Traditional methods rely on linguistic cues such as causal connectives (e.g., “cause”, “because”, and “lead to”) and grammatical patterns to identify causal pairs [104]. A later work utilizes the power of statistical machine learning and deep learning to tackle this task in a supervised learning setting [105]. As LLMs show promising potential with reasoning capacities as introduced in Section 4.1, many works use LLMs as a query tool to determine the edge direction between two given variables. For example, Kıcıman et al. [44] show that LLMs can achieve competitive performance in determining such pairwise causal relationships with accuracies up to $97\%$. Analyses in the medical domain [71, 4, 5] exhibit similar observations. However, other studies highlight LLMs’ limitations of such pairwise causal relationships. For example, sensitivity to prompt design leads to inconsistent results [61]; pairwise judgments can lead to cycles in the full causal graph [92]; pairwise judgments require large computational cost when applying to a large-scale dataset, $N$ variables would require $\binom{N}{2}$ prompts [7]; LLMs still provide false information despite achieving strong results in most cases [61, 91]. Long et al. [61] propose strategies for amending LLMs’ output based on consistency properties in causal inference [61].
To alleviate the impact of erroneous causal information from LLMs, previous works have integrated LLMs with traditional causal discovery methods. Causal discovery or causal structure learning is the task of recovering causal graphs from observational data whenever possible [108]. Traditional causal discovery methods mainly include constrained-based methods that exploit a sequence of statistical tests and score-based methods that structure around the maximization of the fitness of a graph through a space of possible graphs. Vashishtha et al. propose two algorithms that combine LLMs with causal discovery methods: the first uses causal order from an LLM to orient the undirected edges outputted by a constraint-based algorithm and the second utilizes the LLM causal order as a prior for a score-based algorithm [92]. Ban et al. incorporate LLM-driven causal statements as qualitative ancestral constraints in the Bayesian network structure to guide data-driven algorithms [8], which benefits smaller-scale problems, but encounters difficulties with larger datasets due to inaccuracies in the LLM-derived constraints. They then propose an iterative framework that employs LLMs to validate the accuracy of edges in the learned causal graph and fine-tune the causal discovery process based on LLM feedback [7].
6 Future Directions
Theoretical understanding of LLM’s reasoning capacity.
Causal inference methods offer a promising avenue for gaining deeper insights into the reasoning capacity of Large Language Models (LLMs). A potential approach involves employing treatment effect estimation techniques to assess their performance on specific tasks. The gold standard for treatment effect estimation is comparing potential outcomes across various treatments in a controlled experiment. However, the practical challenges of conducting such experiments in real-world scenarios make the interactive nature of LLMs an ideal candidate for investigation. Leveraging their inherent interactivity, researchers could explore quasi-experimental settings, utilizing the natural variations in responses to infer causal relationships. It is crucial to acknowledge the complexities associated with interpreting LLMs, potential biases in training data, and the intricate nature of language understanding tasks. Additionally, considering ethical implications and addressing biases in both training and evaluation data is essential when delving into the reasoning capacity of LLMs using causal inference methods.
Efficient training and inference in LLMs.
As model scales and training data continue to expand, the process of knowledge updating and inference in Large Language Models (LLMs) becomes increasingly resource-intensive. Consequently, it becomes imperative for us to devise methods that can efficiently and judiciously update the knowledge base of pre-trained models. In this context, causal inference methods can play a crucial role by offering guidance in quantifying efficiency. By establishing causal relationships between methods of interest and the existing knowledge collection, these methods can help assess the impact of different updating strategies. This approach not only addresses the resource challenges associated with knowledge updating but also contributes to a more nuanced understanding of the evolving knowledge landscape in LLMs.
LLM-based counterfactual estimation and augmentation.
As a general expert, Large Language Models (LLMs) can significantly contribute to overcoming the current limitations of causal inference methods. A common assumption in many causal methods, as indicated in 5.3, is the existence of corresponding data points for every treatment. However, this assumption often proves untrue, particularly when dealing with imbalanced minority data that may not support meaningful learning. LLMs, functioning as versatile experts, have the potential to address this challenge by aiding in data augmentation for minority data. Through their comprehensive understanding of language and context, LLMs can enhance the availability of diverse data points, facilitating more robust and effective causal inference in situations where traditional methods may struggle due to data imbalances. Similarly, many methods operate under the assumption of unconfoundedness (5.2) within the potential outcome framework, a condition considered quite strong. Historically, this assumption has been accepted due to a lack of domain knowledge regarding the underlying causal graph or identification of potential confounders. However, with the advent of Large Language Models (LLMs), there is an opportunity to alleviate this stringent limitation. LLMs can act as general experts, offering valuable information about potential causal graphs and knowledge. This transformative capability of LLMs opens avenues to enhance our understanding of causal relationships, addressing a historical challenge where unconfoundedness assumptions were made due to limited domain knowledge.
7 Conclusion
At its core, a large language model (LLM) is like a vast library of knowledge. One of the ongoing challenges is figuring out how to extract and use this knowledge effectively. The key to improving LLMs lies in enhancing their ability to understand cause and effect – essentially, how things are connected. Causal reasoning is crucial for making LLMs smarter. Looking at it from a causal inference perspective, we find a valuable framework that helps boost the effectiveness of LLMs. Meanwhile, as keepers of human knowledge, LLMs can even help overcome limitations in causal inference by providing broad expertise that goes beyond existing constraints, reshaping our understanding in this important area and bringing new vitality to this area.
In this survey, we offer a thorough examination of the current landscape where large language models (LLM) intersect with causal inference. We delve into how causal inference contributes to LLM, enhancing aspects such as reasoning, fairness, and safety, along with the explainability of LLM. Additionally, we explore how LLM, in turn, broadens the horizons of causal inference. Across these categories, our survey provides in-depth discussions, comparisons, and concise summaries of the methods scrutinized, offering a comprehensive overview of the current state of research at this intersection. The available benchmark datasets and open-source codes of those methods are also listed.
The examination of the current advancements in causal inference and large language models serves a dual purpose. Firstly, it enhances our comprehension of how these two fields mutually benefit from each other. Secondly, it catalyzes the emergence of novel questions, propelling us closer to achieving Artificial General Intelligence. Moreover, this exploration has the potential to extend into diverse realms and find applications in real-world scenarios, showcasing the far-reaching implications of the synergy between causal inference and LLM.

### v2

@@
 Abstract
-Causal inference has shown potential in enhancing the predictive accuracy, fairness, robustness, and explainability of Natural Language Processing (NLP) models by capturing causal relationships among variables.
-The emergence of generative Large Language Models (LLMs) has significantly impacted various NLP domains, particularly through their advanced reasoning capabilities.
-This survey focuses on evaluating and improving LLMs from a causal view in the following areas: understanding and improving the LLMs’ reasoning capacity, addressing fairness and safety issues in LLMs, complementing LLMs with explanations, and handling multimodality.
-Meanwhile, LLMs’ strong reasoning capacities can in turn contribute to the field of causal inference by aiding causal relationship discovery and causal effect estimations.
-This review explores the interplay between causal inference frameworks and LLMs from both perspectives, emphasizing their collective potential to further the development of more advanced and equitable artificial intelligence systems.
+Causal inference has demonstrated significant potential to enhance Natural Language Processing (NLP) models in areas such as predictive accuracy, fairness, robustness, and explainability by capturing causal relationships among variables.
+The rise of generative Large Language Models (LLMs) has greatly impacted various language processing tasks.
+This survey focuses on research that evaluates or improves LLMs from a causal view in the following areas: reasoning capacity, fairness and safety issues, explainability, and handling multimodality.
+Meanwhile, LLMs can assist in causal inference tasks, such as causal relationship discovery and causal effect estimation, by leveraging their generation ability and knowledge learned during pre-training.
+This review explores the interplay between causal inference frameworks and LLMs from both perspectives, emphasizing their collective potential to further the development of more advanced and robust artificial intelligence systems.
+Large Language Models and Causal Inference in Collaboration: A Comprehensive Survey
+Xiaoyu Liu1,*, Paiheng Xu1,*, Junda Wu2, Jiaxin Yuan1, Yifan Yang1, Yuhang Zhou1, Fuxiao Liu1, Tianrui Guan 1, Haoliang Wang3, Tong Yu3, Julian McAuley2, Wei Ai1, Furong Huang1 1University of Maryland, College Park, 2University of California San Diego, 3Adobe Research Correspondence: furongh@umd.edu.
+* denotes equal contribution.
 1 Introduction
 Recently Large Language Models (LLMs) have showcased remarkable versatility across a spectrum of critical tasks.
-An LLM is adept at tasks such as copywriting, enhancing original sentences with their distinct style and voice, responding to knowledge base queries, generating code, solving mathematical problems, and performing classification or generation tasks tailored to user requirements.
-Moreover, there has been a recent expansion into multi-modal variants, such as Large Visual Language Models (LVLMs) or Large Multi-modal Language Models, which broaden their input/output capabilities to encompass various modalities.
+Moreover, there has been a recent expansion into multi-modal variants, such as Large Vision Language Models (VLMs) or Multi-modal Large Language Models (MLLMs), which broaden their input/output capabilities to encompass various modalities.
 This evolution has significantly enhanced both the potential and range of applications of these models.
-In this survey, our primary focus is on Transformer-based Large Language Models (LLMs).
-The capability of LLMs is fundamentally rooted in their inference abilities, which dictates their proficiency in comprehending, processing, and providing solutions to various inquiries, as well as their adaptability to societally impactful domains.
+In this survey, our primary focus is on Transformer-based LLMs and LVLMs.
+The capability of LLMs is fundamentally rooted in their inference abilities, which dictates their proficiency in comprehending, processing, and providing solutions to various inquiries, as well as their adaptability to societally impactful domains Zhao et al.
+(2023d).
 Consequently, extensive research efforts have been dedicated to measuring and enhancing these capabilities, ranging from assessing the reasoning abilities of LLMs to scrutinizing their decision-making processes and addressing challenges such as concept alignment across different modalities and mitigating hallucination.
-In addition, since LLMs are trained on extensive human knowledge with billions of parameters, they sometimes face challenges in appropriately prioritizing or downplaying what they have learned in different scenarios.
-This can lead to issues such as domain shift, where the model’s performance degrades on data that differ from the training set, and long-tail bias, where infrequent examples are not handled as effectively.
-In many instances, language tasks require not only predicting or generating text based on patterns in the data but also understanding the underlying causal mechanisms driving these patterns.
-Causal inference has shown great potential in improving predictive accuracy, fairness, robustness, and explainability of Natural Language Processing (NLP) models [23].
-With the advent of generative LLMs, a significant transformation has occurred across various NLP fields, attracting increased research interest in applying causal inference to address LLM-related challenges and augment their functionality.
-Such transformation also motivates this survey to outline causal methodologies and their implementation in LLMs, emphasizing their role in enriching our comprehension and application of language models.
+Meanwhile, causal inference has shown great potential in improving predictive accuracy, fairness, robustness, and explainability of Natural Language Processing (NLP) models Feder et al.
+(2022), With LLMs revolutionizing various language processing tasks, there is a growing trend in applying causal inference to address LLM-related challenges and enhance their functionality.
+This survey outlines causal methodologies and their implementation in LLMs, emphasizing their role in enriching our comprehension and application of language models.
+Moreover, this survey also aims to explore how LLMs can help with the causal inference framework.
+Causal inference is formally defined as an intellectual discipline that considers the assumptions, study designs, and estimation strategies that allow researchers to draw causal conclusions based on data Pearl (2009).
+Causal inference has three main origins: potential outcomes, graphs, and structural equations, each serving unique purposes Zeng and Wang (2022).
+In this survey, we mainly discuss Pearl’s formulation of causal graphs Pearl (1998), which formalized causal graphical models for presenting conditional independence among random variables using directed acyclic graphs (DAGs).
+We summarize how LLMs can help causal inference in its two important components, i.e., causal relationship discovery and treatment effect estimation.
+Estimating causal effects between variables requires assumptions about their relationships with other variables, traditionally provided by experts LLMs, leveraging pre-trained knowledge, can assist in identifying these relationships and enhance causal discovery methods Zanga et al.
+(2022) for more reliable outcomes.
+Additionally, estimating treatment effects is often hindered by the absence of counterfactual data.
+By utilizing LLMs’ strong generative abilities, researchers have developed various ways to generate high-quality counterfactuals to enable treatment effect estimation.
+Figure 1 shows an overview of selected topics about the interplay between causal inference frameworks and LLMs from both perspectives.
 {forest}
-for tree= draw=red, align=center, anchor=west, child anchor=west, grow’=0, rounded corners, font= [Causal and Large Language Models (LLMs), rotate=90, anchor=south, text width=5.5cm [Causal for LLMs, [Reasoning Capacity [Model Understanding [[109, 79, 43, 47, 1, 88, 116], fill=yellow!20] ] [Commonsense Reasoning [[27, 98, 114, 123, 62, 17], fill=yellow!20] ] [Counterfactual Reasoning [[11, 70, 50, 60], fill=yellow!20] ] ] [Fairness and Bias [[68, 2, 82, 22, 124, 95, 65, 37], fill=yellow!20] ] [Safety [[15, 9, 122, 118, 69, 56], fill=yellow!20] ] [Explanation [[35, 10, 33, 119, 55], fill=yellow!20] ] [Multi-modality [[74, 45, 49, 85, 86, 121], fill=yellow!20] ] [Benchmark [[103, 115, 36, 107, 41, 72, 40, 85, 49, 18, 27], fill=yellow!20] ] ] [LLMs for Causal [Treatment Effect Estimation [[19, 24, 113], fill=yellow!20] ] [Causal Relationships Discovery [[44, 71, 4, 5, 61, 92, 7, 91, 8, 7], fill=yellow!20] ] ] ]
-Figure 1: Structure Overview
-Meanwhile, this survey also aims to explore how LLMs can help with the causal inference framework.
-Causal inference is formally defined as an intellectual discipline that considers the assumptions, study designs, and estimation strategies that allow researchers to draw causal conclusions based on data [75].
-Causal inference has three main origins: potential outcomes, graphs, and structural equations, each serving unique purposes [110].
-The potential outcomes framework [75] focuses on estimating causal effects through statistical inference and treatment comparisons.
-Graphical models, meanwhile, excel in mapping out causal pathways and visualizing relationships, with nodes representing variables and edges denoting directional influences.
-In this survey, we mainly discuss Pearl’s formulation of causal graphs [76], which formalized causal graphical models for presenting conditional independence relations among random variables using directed acyclic graphs (DAGs).
+for tree= draw=red, align=center, anchor=west, child anchor=west, grow’=0, rounded corners, font= [Causal and Large Language Models (LLMs), rotate=90, anchor=south, text width=5.5cm [Causal for LLMs, [Reasoning Capacity [Model Understanding [Zečević et al.
+(2023); Romanou et al.
+(2023); Kim et al.
+(2023); Li and Li (2023)
+Abdali et al. (2023); Tang et al. (2023); Zhang et al. (2023c); Chen et al. (2024b), fill=yellow!20] ] [Commonsense Reasoning [Gao et al. (2023); Willig et al. (2022); Zhang et al. (2022)
+Zheng et al.
+(2023); Lu et al.
+(2022); Chen et al.
+(2023a), fill=yellow!20] ] [Counterfactual Reasoning [Betti et al.
+(2023); Miao et al.
+(2023); Li et al.
+(2023c); Liu et al.
+(2023d), fill=yellow!20] ] ] [Fairness and Bias [Meade et al.
+(2021); Achiam et al.
+(2023); Stanovsky et al.
+(2019); Ding et al.
+(2022); Zhou et al.
+(2023a); Wang et al.
+(2023)
+Madhavan et al.
+(2023); Jenny et al.
+(2023); Xia et al.
+(2024); Wu et al.
+(2024a); Gallegos et al.
+(2024), fill=yellow!20] ] [Safety [Cao et al.
+(2022); Bao et al.
+(2021); Zhao et al.
+(2023e, 2022); Meng et al.
+(2021); Liu et al.
+(2023a), fill=yellow!20] ] [Explanation [Hou et al.
+(2023); Belrose et al.
+(2023); Gurnee et al.
+(2023); Zhao et al.
+(2023a); Liu et al.
+(2023b); Vig et al.
+(2020), fill=yellow!20] ] [Multi-modality [Pawlowski et al.
+(2023); Ko et al.
+(2023); Li et al.
+(2023b); Su et al.
+(2023); Tai et al.
+(2023); Zhao et al.
+(2023c); Chen et al.
+(2024a), fill=yellow!20] ] [Benchmark [Xie et al.
+(2023); Zhang et al.
+(2023b); Huang et al.
+(2023); Yu et al.
+(2023); Jin et al.
+(2023a); Nie et al.
+(2023)
+Jin et al.
+(2023b); Su et al.
+(2023); Li et al.
+(2023b); Chen et al.
+(2023b); Gao et al.
+(2023); Chen et al.
+(2024b, a), fill=yellow!20] ] ] [LLMs for Causal [Treatment Effect Estimation [Chen et al.
+(2023c); Feder et al.
+(2023); Zhang et al.
+(2023a), fill=yellow!20] ] [Causal Relationships Discovery [Kıcıman et al.
+(2023); Naik et al.
+(2023); Antonucci et al.
+(2023); Long et al.
+(2023); Vashishtha et al.
+(2023)
+Arsenyan and Shahnazaryan (2023); Ban et al. (2023a); Tu et al. (2023); Ban et al. (2023b, a)
+Joshi et al. (2024); Feng et al. (2024); Li et al. (2024); Khatibi et al. (2024); Wan et al. (2024), fill=yellow!20] ] ] ]
+Figure 1: An overview of the interplay between causal inference frameworks and LLMs.
+2 Causal Inference and Large Language Models
+In this section, we provide a brief introduction to both large language models (LLMs) and causal inference, laying the groundwork for understanding their interaction.
+We begin by outlining the core principles and recent advancements in LLMs, which have reshaped natural language processing by enabling sophisticated language generation and comprehension.
+Following this, we introduce the fundamental concepts of causal inference, emphasizing its role in discerning cause-and-effect relationships within data.
+By establishing a foundational understanding of these two fields, we aim to clarify how their integration can lead to significant advancements, enhancing both the reasoning capabilities of LLMs and the development of causal inference methodologies.
+2.1 Evolution of the Large Language Models
+Large Language Models (LLMs) have transformed the way we interact with and process language, opening up new possibilities for natural language understanding, generation, and communication Zhao et al.
+(2023d).
+In this paper, we mainly focus on Transformer Vaswani et al.
+(2017) based LLMs, and we provide an overview of their recent progress in this section.
+The introduction of the highly parallelizable Transformer architecture Vaswani et al.
+(2017), which leverages self-attention mechanisms, has led to the development of a range of Pre-trained Language Models (PLMs) with varying architectures and pre-training strategies, e.g., BERT Devlin et al.
+(2019), GPT2 Radford et al.
+(2019), and so on Radford et al.
+(2018); Liu et al.
+(2019); Lewis et al.
+(2020).
+These models largely improved performances for NLP tasks with a learning paradigm of general-purpose pre-training and task-specific fine-tuning.
+Researchers then find that scaling the size of the model or pre-training data often leads to further improved model capacity on downstream tasks Kaplan et al.
+(2020); Wei et al.
+(2022a); Schaeffer et al.
+(2024).
+These larger-scaled PLMs are referred to as large language models (LLMs) Brown et al.
+(2020); Chowdhery et al.
+(2022); Team et al.
+(2023); Achiam et al.
+(2023); Bai et al.
+(2023); Touvron et al.
+(2023); Dubey et al.
+(2024).
+These LLMs demonstrated strong capacities for language processing and solving complex tasks through text generation.
+We briefly introduce three common strategies that are used to further enhance their performance: In-context Learning (ICL) Brown et al.
+(2020), which allows LLMs with natural language instruction and/or several task demonstrations (i.e., input-output pairs); Chain-of-Through (CoT) prompting Wei et al.
+(2024), which includes intermediate reasoning steps into prompts to guide the model toward the final answer; and Instruction tuning Sanh et al.
+(2024); Ouyang et al.
+(2022); Thoppilan et al.
+(2022), which refers to fine-tuning the LLMs with a mixture of multi-task datasets formatted as natural language instructions.
+Additionally, researchers have extended these LLMs to handle images as input Team et al.
+(2023); Achiam et al.
+(2023), leading to the development of Large Vision Language Models (VLMs) or Multi-modal Large Language Models (MLLMs).
+In this survey, we primarily focus on works that adopt a causal perspective to improve LLMs in their generative capacities.
+2.2 Introduction of Causal Inference
+In this section, we present the background knowledge of causal inference, including task descriptions, basic concepts and notations, and general solutions. More details can be found in Appendix A.
+Causal inference aims to estimate the causal relationship among variables.
+The variables of interest are referred to as treatment, naturally, the effects of treatments are referred to as treatment effects.
+Ideally, the treatment effect can be measured as follows: applying different treatments to the same cohort, and then the difference in the effect is the treatment effect.
+However, it is impracticable for perfectly controlled experiments in many cases, requiring estimation of treatment effect from observational data.
+One of the most influential frameworks in identifying and quantifying causal effects in observational data is the potential outcomes framework Rubin (1974).
+The potential outcomes approach associates causality with manipulation applied to units, and compares causal effects of different treatments via their corresponding potential outcomes.
+Definition 2.1 (Binary Average Treatment Effect(ATE)).
+Suppose we want to measure the treatment effect of a treatment $T=1$. Then the average treatment effect is defined as:
+$\mathbb{E}[Y(T=1)-Y(T=0)]$ (1)
+where $Y(T=1)$ and $Y(T=0)$ denote the potential treated and control outcome of the whole population respectively.
+The potential outcome framework is powerful in recovering the effect of causes.
+In a potential outcome framework, causal effects are answered by specific manipulation of treatments.
+However, when it comes to identifying the causal pathway or visualizing causal networks, the potential outcome model has its limitations.
+In the front of the challenge, causal graphical models utilize directed edges to represent causalities and encode conditional independence among variables in the graphs.
+One of the most widely-spread formulations is the Structural Equation Model Wright (1934); Pearl (1998), where linear structural equation models are used to present causal relationships by directed edges, which differentiate correlation from causation when the graph structure is given.
+The linearity assumption was later relaxed by Pearl (1998) and it formalized causal graphical models for presenting causal relations using Directed Acyclic Graphs (DAGs).
+Specifically, consider the random variable $\mathbf{X}\in\mathcal{R}^{D\times N}=[X_{1},X_{2},...,X_{N}]$, where $D$ represents the dimensionality of each variable, and $N$ is the number of variables, the linear SEM consists of a set of equations of the form:
+$X_{i}=\beta_{0i}+\sum_{j\in pa(X_{i})}\beta_{ji}X_{j}+\epsilon_{i},\quad i=1,2,3,...,N$ (2)
+where $pa(X_{i})$ denotes the set of variables that are direct parents of $X_{i}$.
+$\epsilon_{1},\epsilon_{2},...,\epsilon_{N}$ are mutually independent noise terms with zero mean, $\beta_{ji}$ are coefficients that quantify the causal effect of $X_{j}$ on $X_{i}$.
+While the non-parametric SEM takes the form:
+$X_{i}=f_{i}(\mathbf{X}_{pa(i)},\epsilon_{i}),\quad i=1,2,3,...,N$ (3)
+The random variables $\mathbf{X}$ that satisfies the model structure of the form in Equation 6 or Equation 7 can be represented by a directed acyclic graph (DAG) $G=(V,E)$, where $V$ is the set of associated vertices, each corresponding to one of a variable of interest $X_{i}$, and $E$ is the corresponding edge set.
+With pre-specified DAG and assumptions on the latent variables, the coefficients between the latent variables are identifiable Kuroki and Pearl (2014).
+Next, we show a comprehensive survey of how existing causal frameworks help challenges in LLMs.
+3 Causal Inference for Large Language Models
+LLMs can significantly benefit from causal inference as it enhances their ability to understand and reason about cause-and-effect relationships within data.
+In this section, we review how LLMs can benefit from a causal lens in various capacities.
+We show an overview in Figure 1.
+3.1 Reasoning Capacity
+3.1.1 Model Understanding
+LLMs have demonstrated many emerging abilities in language generation and certain reasoning tasks Bubeck et al.
+(2023); Kıcıman et al.
+(2023).
+As the reasoning process is often associated with causal factors, it is logical to first understand and evaluate LLMs’ reasoning ability from a causal lens.
+Zečević et al.
+(2023) argued LLMs are not causal and hypothesized that LLMs are simply trained on the data, in which causal knowledge is embedded.
+Thus, in the inference stage, the LLMs can directly recite the causal knowledge without understanding the true causality in the context.
+Similar behaviors are exhibited in a Causal Reasoning Assessment Benchmark, CRAB Romanou et al.
+(2023).
+Kim et al.
+(2023) examined LLMs’ abilities to understand the causalities of both scientific papers and newspapers.
+The results show that ChatGPT performs worse than a fine-tuned BERT model in determining the causality or correlation of given statements.
+Abdali et al.
+(2023) show the effectiveness of applying LLMs to diagnose the cause of issues from Microsoft Windows Feedback Hub.
+Li and Li (2023) showed that LLMs can identify dynamical (spatio-temporal) effects.
+However, how to infer the relationship and interactions of them is still challenging for LLMs, which are more emphasized as causal structures in causal inference.
+Another important line of work is to understand LLMs’ hallucination and faithfulness in knowledge reasoning by considering causal effects.
+Tang et al.
+(2023) proposed a multi-agent system, CaCo-CoT, where some LLMs are reasoners and others are evaluators.
+Reasoners try to provide causal solutions, while evaluators try to challenge the reasoners with counterfactual candidates.
+With the cooperative reasoning framework, CaCo-CoT helps to improve causal-consistency.
+Zhang et al.
+(2023c) identified the potential knowledge bias pretrained in the LLMs as the confounder which causes incorrect answers and hallucinations.
+They proposed a CoT framework to generate sub-questions necessary to answer a question and require humans to provide the correct answers.
+Moreover, Chen et al.
+(2024b) introduced CaLM, a comprehensive benchmark that systematically assesses LLMs’ causal reasoning across diverse tasks, adaptations, and error types.
+3.1.2 Commonsense Reasoning
+Commonsense reasoning involves the ability to apply everyday knowledge and intuitive understandings of the world to make decisions or draw conclusions, which is vital for LLMs’ contextual understanding and human-like interactions Davis and Marcus (2015); Storks et al.
+(2019).
+This section briefly summarizes the commonsense reasoning ability of LLMs under various settings Gao et al.
+(2023); Willig et al.
+(2022) and the employment of causally motivated methods in improving commonsense causality reasoning Zhang et al.
+(2022); Zheng et al.
+(2023); Lu et al.
+(2022); Chen et al.
+(2023a); Zhao et al.
+(2023b).
+The reasoning ability of LLMs is limited but it could generate good causal explanations Gao et al.
+(2023).
+This viewpoint has been validated through ChatGPT on event causality identification, causal discovery, and causal explanation generation.
+Willig et al.
+(2022) showed a similar observation by evaluating LLMs on causal question answering, where LLMs arrived at their answers through memorization instead of reasoning.
+To improve the commonsense causality reasoning of LLMs that identify causes from effects in natural language, ROCK Zhang et al.
+(2022) balances confounding effects using temporal propensities through an estimation of the average treatment effect.
+While ROCK adopts a potential outcome framework, Chen et al.
+(2023a) uses a conversation cognitive model based on intuition theories and transforms intuitive reasoning into a structural causal model.
+Other than facilitating the reasoning ability of LLMs directly, Zheng et al.
+(2023) use causal inference to preserve commonsense knowledge from pre-trained language models for fine-tuning to prevent catastrophic forgetting; Lu et al.
+(2022) focus on improving LLM’s ability in generalized procedural planning with commonsense-infused prompts.
+For procedural planning tasks, Lu et al.
+(2022) proposed to learn causeeffect relations among complex goals and stepwise tasks, and reduced spurious correlation among them via front door adjustment.
+3.1.3 Counterfactual Reasoning
+LLMs can generate counterfactuals as data augmentations for small language models.
+Given a text $x$ and a black-box classifier $B$, the counterfactual text $\tilde{x}$ of text $x$ should satisfy: Betti et al.
+(2023); Miao et al.
+(2023): (1) $\tilde{x}$ has a different class than $x$, $B(x)\neq B(\tilde{x})$; (2) $x$ and $\tilde{x}$ differ only by minimal lexical changes; (3) $\tilde{x}$ is a feasible text and the commonsense constraint is satisfied.
+To understand LLM’s ability to generate counterfactuals, Li et al.
+(2023c) examined the following tasks: sentiment analysis , natural language inference (NLI) , named entity recognition (NER) , and relation extraction (RE) For simple tasks like sentiment analysis and NLI, data augmented via LLMs can mitigate potential spurious correlations.
+For more complicated tasks like RE, LLMs may generate low-quality counterfactuals.
+Liu et al.
+(2023d) evaluated abductive reasoning and counterfactual reasoning abilities and found code large language models (Code-LLMs) achieved better results compared to text models.
+Many research also showed different ways of incorporating counterfactuals in domain-specific tasks.
+Miao et al.
+(2023) claimed that existing methods for generating counterfactuals for RE face two challenges: identifying causal terms correctly and ignoring the commonsense constraint.
+To amend this, they proposed an intervention-based strategy to generate commonsense counterfactuals for stable relation extraction.
+Similarly, Oba et al.
+(2023) and Sen et al.
+(2023) used counterfactual generation to address issues in gender bias and harmful language detection, respectively.
+3.2 Fairness and Bias
+Fairness and bias are pivotal factors in deploying language models effectively and ethically.
+Bias is common in pretrained language models as they capture and potentially amplify undesired social stereotypes and biases Meade et al.
+(2021); Achiam et al.
+(2023); Zhou et al.
+(2024); Wu et al.
+(2024b); Wang et al.
+(2024); Gallegos et al.
+(2024).
+An example of bias in language models includes gender associations with specific professions, such as male firefighters and female nurses Stanovsky et al.
+(2019).
+Causality-based methodologies offer a promising approach for mitigating biases in language models by discerning the origins of bias through a causal perspective.
+Bias mitigation is then followed by eliminating the unwanted spurious correlation between generative factors through different types of causal intervention or causal invariant learning.
+Ding et al.
+(2022) introduced a proxy variable related to gender bias in the causal graph, and used two different ways to eliminate the potential proxy bias and unresolved bias under the linear structural equation model.
+Zhou et al.
+(2023a) believe that the backdoor path between the ground truth label and the non-causal factors is the source of bias, and used the Independent Causal Mechanism principle to mitigate bias.
+Wang et al.
+(2023), from a different angle, eliminated the bias by performing a do operation on the intermediate variables for both white-box and black-box LLMs.
+Madhavan et al.
+(2023) considered the tokens generated by generative language models trained with causal language modeling objectives as a causal graph, and analyzed the bias under this model.
+Xia et al.
+(2024) used a reward model as an instrumental variable to perform causal interventions to reduce biases in LLM outputs by modeling the confounders in pretraining data and input prompts.
+Wu et al.
+(2024a) used external knowledge as an instrumental variable and front-door adjustment to improve the accuracy of CoT reasoning on knowledge-intensive tasks.
+Jenny et al.
+(2023) used Activity Dependency Networks to describe the causality effect between normative variables, such as clarity and authenticity, to structure the cause of bias.
+3.3 Safety
+As researchers have observed the unreliability phenomenon of LLMs in various applications Cao et al.
+(2022); Zhao et al.
+(2023e, 2022); Meng et al.
+(2021); Liu et al.
+(2023a); Xu et al.
+(2024), there is increasing interest in applying causal inference techniques to analyze the causality of the non-robustness of the model and adjust the treatment to resolve the challenges Cao et al.
+(2022); Zhao et al.
+(2022, 2023e).
+LLMs face challenges of unreliability when performing knowledge probing by query LLMs with task-specific prompts Cao et al.
+(2022); Xiao et al.
+(2023), such as using shortcuts to complete the probing and generating different predictions for the semantically equivalent prompt.
+By simply translating text tokens in input prompts to emoji sequences, LLMs generated more severe hallucinations Zhao et al.
+(2023e).
+The main reason is that they rely on spurious correlations to make an inference Zhao et al.
+(2022).
+Training LLMs to learn the causal relationship between input $x$ and output $y$ is an intuitive method of better resisting prompt attacks.
+The randomized smoothing technique Zhai et al.
+(2020); Jia et al.
+(2019) can model the interventional distribution $p(y|do(x))$ by assuming discrete adversarial perturbations as the Gaussian distribution Zhao et al.
+(2022).
+The method of smoothing in the latent semantic space is more robust against known attacks such as word substitutions, paraphrasing, and token position change Zhao et al.
+(2022).
+3.4 Explainability
+Explainability in LLMs refers to the capacity to elucidate how these models arrive at their conclusions, enhancing transparency and trustworthiness in AI decision-making processes Guidotti et al.
+(2018); Lipton (2018).
+Many works have tried to understand the inner workings of LLMs Hou et al.
+(2023); Belrose et al.
+(2023); Gurnee et al.
+(2023); Zhao et al.
+(2023a); Liu et al.
+(2023b); Li et al.
+(2023d).
+We summarize research efforts that probe the causal mechanism in LLMs from the following three directions: intervening the inputs or prompts, intervening inner components of LLMs, and abstracting the working mechanism into a causal graph.
+Inputs or Prompt Intervention
+Input intervention, as a data-centric method, is to create counterfactual input text by changing the treated feature in the text, and then observe the model behaviors on the original and counterfactual texts.
+LLMs can first identify the features in input texts causally associated with the predictions and are capable of changing the identified features to create the counterfactual texts Bhattacharjee et al.
+(2023); Gat et al.
+(2023).
+These counterfactual texts can be utilized to investigate the causality of the LLM and can serve as a training dataset to learn a matching model, where the matched counterfactual pairs have similar embeddings Gat et al.
+(2023).
+Various works have developed different prompting methods and found whether the prompting methods are causally associated with the final output of LLMs Schick and Schütze (2020); Zhou et al.
+(2023b); Wei et al.
+(2022b).
+However, the causal effect of prompting methods, such as chain-of-thought (CoT), on the final output is ambiguous.
+Prompt intervention, which alters only one particular aspect of prompts, is proposed to understand the contributions of each component of prompts on model behavior Madaan et al.
+(2023); Ji et al.
+(2023); Tan (2023).
+These findings suggest that LLMs rely on the causal model suggested by their CoTs to a high extent, but LLMs also learn spurious correlations such as sentence length to generate responses.
+Inner Component Intervention
+Input or prompt interventions help understand model behavior, but inner component intervention is needed to uncover the information cascade within LLMs.
+Existing works mainly focus on the essential components in SoTA LLMs: attention mechanism and MLP.
+Stolfo et al.
+(2023) exchanged the activation values in MLP and attention layers of different inputs to probe the function of MLPs and attention mechanism.
+While Stolfo et al.
+(2023) focuses only on math word problems with four fundamental arithmetic operators, it is an interesting direction to generalize the component intervention to other applications.
+Vig et al.
+(2020) used causal mediation analysis to interpret which components of neural models are causally implicated in their behaviors and applies this approach to uncover how gender bias is mediated in GPT2.
+Causal Graph Abstraction
+An intuitive way to characterize causality within LLMs is to abstract the working mechanism of LLMs into a causal graph.
+Boundless Distributed Alignment Search (DAS) Wu et al.
+(2023), by replacing brute-force searching the original DAS Geiger et al.
+(2023) with learnable parameters, has been effective on the Alpaca model Taori et al.
+(2023).
+Given four pre-defined causal models, Boundless DAS extracts two of them as the accurate hypotheses as the abstracted causal graph of the Alpaca model.
+However, the Boundless DAS method is restricted by the given causal hypothesis, and the future direction can explore how to abstract the causal graph in LLMs without prior causal graphs.
+3.5 Multi-modality
+Large vision-language models have been popular in many applications Maaz et al.
+(2023); Li et al.
+(2023a); Liu et al.
+(2023c); Guan et al.
+(2024).
+How to conduct causal reasoning on both images and texts can be crucial in correctly answer multimodal questions Niu et al.
+(2021).
+Pawlowski et al.
+(2023) examined LLMs’ causal reasoning abilities and showed that the causal knowledge in the language models can be too strong a prior which often causes the model to neglect visual evidence.
+Si et al.
+(2022) highlighted how VQA models tend to rely on various shortcuts beyond language priors.
+Ko et al.
+(2023) proposed to alleviate the problem by adding self-consistent generation prediction by checking predictions based on different parts of the input, in which the three inputs V, Q, and A are individually predicted based on the other two inputs.
+Additionally, Li et al.
+(2023b) proposed an image generation framework with causal reasoning and created a novel VQA dataset whose questions require causal explanations.
+Chen et al.
+(2024a) curated a multi-hop VQA dataset, assessed unimodal biases in LVLMs, and proposed a causality-enhanced agent framework to mitigate them.
+Another important question is to understand the spatial-temporal causal relationship of the visual elements within the images and videos.
+Su et al.
+(2023) proposed CaKe-LM to use pretrained causal knowledge for video understanding while Tai et al.
+(2023); Zhao et al.
+(2023c) designed specific structure and prompting method to capture and interpret the underlying causal relationship for VQA.
+3.6 Evaluation and Benchmark
+In this section, we list existing LLM benchmarks built from a causal perspective.
+Reference MU CR CF FD Text MM
+ECHo Xie et al. (2023)
+CREPE Zhang et al. (2023b)
+CLOMO Huang et al. (2023)
+IfQA Yu et al. (2023)
+Cladder Jin et al. (2023a)
+MoCa Nie et al. (2023)
+Jin et al. (2023b)
+CVidQA Su et al. (2023)
+VQAI Li et al. (2023b)
+Chen et al. (2023b)
+Gao et al. (2023)
+CRAB Romanou et al. (2023)
+HELM Liang et al. (2022)
+Fair-Prism Fleisig et al. (2023)
+Biasasker Wan et al. (2023)
+CaLM Chen et al. (2024b)
+MORE Chen et al. (2024a)
+Table 1: Existing evaluation benchmarks.
+Based on the evaluation tasks, we categorize the benchmarks into three categories: Model Understanding (MU), Commonsense Reasoning (CR), Counterfactual Reasoning (CF) and Fairness/Debias (FD).
+Based on the modalities of the data samples, we identify the benchmarks with only textual inputs (Text) and those with multimodal inputs (MM).
+The benchmarks in model understanding (MU) focus on evaluating and understanding LLMs’ causal reasoning abilities in both natural language Jin et al.
+(2023a); Nie et al.
+(2023); Jin et al.
+(2023b); Gao et al.
+(2023); Romanou et al.
+(2023); Chen et al.
+(2024b) and vison-language Su et al.
+(2023); Li et al.
+(2023b); Xie et al.
+(2023); Chen et al.
+(2024a).
+In addition, some benchmarks Nie et al.
+(2023); Gao et al.
+(2023) also provide model understanding in comparison with human causal reasoning and moral judgments.
+Commonsense reasoning benchmarks (CR) evaluate LLMs on tasks that require extensive commonsense knowledge for both textual-only context Zhang et al.
+(2023b); Jin et al.
+(2023a) and multimodal context Xie et al.
+(2023).
+Contexts with commonsensical and anti-commonsensical are constructed in Jin et al.
+(2023a), to further investigate whether LLMs use averaged-out causal reasoning.
+Evaluating LLMs’ counterfactual reasoning (CF) abilities is essential in enabling explainable model reasoning and calibration of the generated rationales.
+Huang et al.
+(2023) introduce a specific task and benchmark for assessing LLMs’ logical counterfactual thinking.
+Yu et al.
+(2023) contribute a novel dataset to challenge LLMs in counterfactual reasoning in an open-domain QA context.
+Chen et al.
+(2023b) investigate the ability of LLMs to provide explanations that aid in understanding their reasoning process, particularly in the context of counterfactual scenarios.
+The fairness and bias (FD) evaluations are particularly in addressing biases, fairness, and the overall transparency of language models.
+HELM Liang et al.
+(2022) is a comprehensive evaluation benchmark including previously neglected areas for fairness.
+Fair-Prism Fleisig et al.
+(2023) focuses specifically on fairness-related harms in models, which are identified and measured by detailed human annotations.
+Biasasker Wan et al.
+(2023) presents an automated framework to identify and measure social biases by probing the models with specially designed questions.
+4 Large Language Model for Causal Inference
+Causal inference, as a powerful tool for developing large language models (LLMs), greatly benefits from the extensive world knowledge that LLMs provide.
 We summarize how LLMs can help causal inference in its two important components, i.e., causal relationship discovery and treatment effect estimation.
-Determining causal relationships among variables is a fundamental step in a causal inference framework because the estimation of the causal effect of variable $A$ on variable $B$ requires causal assumptions on causal relationships of other variables with $A$ and $B$.
-Traditionally, researchers rely on experts with subject matter knowledge to set the ground for such causal relationships.
-Causal discovery methods [108] provide an alternative by discovering causal graphs from observational data.
-LLMs have demonstrated abilities to determine such causal relationships based on pre-trained knowledge or given text.
-They can also be integrated with causal discovery methods to further increase the reliability of the outcome.
+4.1 Treatment Effect Estimation
 Estimating treatment effects is central to causal inference but is hindered by the absence of counterfactual data in many cases.
-Leveraging the strong counterfactual reasoning abilities of LLMs, researchers have developed various ways to generate high-quality counterfactuals to enable treatment effect estimation.
-The structure of the survey is given in Figure 1.
-We start with an introduction to recent advancements in Large Language Models in Section 2.
-Then we provide an overview of causal inference methods that are used and can be used to improve LLMs in Section 3.
-In the first half of the paper we go through how these methods are used in various problems in the LLM community: Section 4.1 overviews how causal methods are used to measure and improve the reasoning capacity of LLM, Section 4.2 and Section 4.3 focus on the fairness and safety issues while Section 4.4 introduces how the explainability of LLM are approached by causal inference methods.
-We also discuss the extension to construction and development of multi-modality large models in Section 4.5.
-Finally, we list existing work on evaluation and benchmarking LLMs from a causal perspective in Section 4.6.
-In the second half of the survey, we shift to how LLM extends the boundary of causal inference.
-Section 5.1 explains current assumptions, limitations and bottlenecks of causal inference.
-Section 5.3 and Section 5.2 states current works on improving treatment effect estimation and causal discoveries.
-We highlight several future directions in Section 6.
-2 Background of Large Language Models
-Large Language Models (LLMs) have transformed the way we interact with and process language, opening up new possibilities for natural language understanding, generation, and communication [34].
-These models are in a constant state of evolution, continually expanding the limits of what is achievable in language processing and artificial intelligence [73].
-In this paper, we mainly focused on Transformer-based LLMs, and we provide an overview of their recent progress in this session.
-The major breakthrough of Large Language Models (LLM) came in 2017 when the Transformer [93] was introduced.
-In a groundbreaking shift, the Transformer technology emerged, mastering the art of grasping long-term linguistic connections.
-This innovation wasn’t just theoretical—it allowed for simultaneous training on multiple GPUs, paving the way for the creation of significantly larger models [16].
-Then came a pivotal moment in 2018: the unveiling of OpenAI’s GPT-1 [78].
-This was more than just another step in natural language processing (NLP); it was a leap, thanks to its transformer-based architecture.
-Boasting 117 million parameters, GPT-1 wasn’t merely crunching data; it was weaving contextually coherent sentences, showcasing the transformative power of transformers in redefining NLP tasks [3].
-Despite its initial limitations, GPT-1 didn’t just make a mark; it laid the groundwork for a new wave of AI exploration and sparked intense competition in the realm of LLMs.
-In 2020, OpenAI released GPT-3 [20], which was able to generate highly coherent and natural-sounding text [13].
-It was a big deal because it showed just how awesome these huge language models could be at a bunch of different language tasks [26].
-Riding on the high from GPT-3’s success, OpenAI released the next iteration of their language model, GPT-4 [2, 67, 52] with the ability to generate even more coherent and natural-sounding text.
-Following GPT-4’s success, Google wasn’t far behind with their Bard [66].
-Amazon jazzed up Alexa [42] with some cool AI features, Huawei jumped in with their Pangu models [111] and Alibaba proposed their QWEN models [6].
-Then Meta brought out this thing called LLaMA [90], which was all about the first open-source foundation models.
-Compared to LLaMA, LLaMA2 [90] has made more explorations in reinforcement learning from human feedback (RLHF) and developed a chat-oriented version called LLaMA-chat, which generally outperforms existing open-source models across a range of helpfulness and safety benchmarks.
-A large number of researchers have extended LLaMA models by either instruction tuning [57] or continual pretraining.
-Alpaca [89] is the first open instruct-following model fine-tuned based on LLaMA.
-In addition, Vicuna [77] is another popular LLaMA variant, trained upon user-shared conversations collected from ShareGPT [31].
-As diverse versions of LLMs emerge, they encounter common challenges.
-In this survey, we demonstrate that many of these challenges can be effectively addressed through causal methods.
-These methods encompass enhancing the reasoning capacity of LLMs, tackling fairness concerns and mitigating potential biases, ensuring safety, and enhancing the explainability of model outputs, and their extension to multi-modality versions.
-Building upon this progress, there is now a burgeoning interest in expanding the scope of these models to encompass visual data, giving rise to the emergence of Large Visual Language Models (LVLMs).
-These models aim to integrate the understanding of both textual and visual information, opening up new avenues for more comprehensive AI systems capable of interpreting and generating content in multimodal formats.
-One of the most prevalent approaches is inserting visual features as supplementary inputs to LLMs and aligning them with textual features.
-This method has been adapted in several large vision-language models (LVLMs) such as MiniGPT-4 [127], LLaVA [59], Mplug-Owl [106] and so on [55, 57, 117, 96, 106, 58, 51].
-In this survey, we also demonstrate that the causal methods could help concur challenges encountered in existing LVLMs as above.
-3 Brief Introduction of Causal Inference
+Chen et al.
+(2023c) proposed a new method for automatically generating high-quality counterfactual data at scale called DISCO (DIStilled COunterfactual Data).
+In addition,  Feder et al.
+(2023) apply treatment effect estimation to align knowledge for generalization towards different domains.
+Zhang et al.
+(2023a) optimize the treatment effect estimation on unlabeled datasets by performing self-supervised causal learning through LLMs.
+4.2 Causal Relationships Discovery
+Discovering causal relationships between variables is fundamental in causal inference as it enables the identification and estimation of causal effects.
+Many research focus on casual relationship extraction or causality extraction which extracts causal relationships between two variables from text directly.
+Traditional methods rely on linguistic cues such as causal connectives (e.g., “cause”, “because”, and “lead to”) and grammatical patterns to identify causal pairs Xu et al.
+(2020).
+A later work utilizes the power of statistical machine learning and deep learning to tackle this task in a supervised learning setting Yang et al.
+(2022).
+As LLMs show promising potential with reasoning capacities as introduced in Section 3.1, many works use LLMs as a query tool to determine the edge direction between two given variables.
+For example, Kıcıman et al.
+(2023) show that LLMs can achieve competitive performance in determining such pairwise causal relationships with accuracies up to $97\%$.
+Analyses in the medical domain Naik et al.
+(2023); Antonucci et al.
+(2023); Arsenyan and Shahnazaryan (2023) exhibit similar observations.
+However, other studies highlight LLMs’ limitations of such pairwise causal relationships.
+For example, sensitivity to prompt design leads to inconsistent results Long et al.
+(2023); pairwise judgments can lead to cycles in the full causal graph Vashishtha et al.
+(2023); pairwise judgments require large computational cost when applying to a large-scale dataset, $N$ variables would require $\binom{N}{2}$ prompts Ban et al.
+(2023a); LLMs still provide false information despite achieving strong results in many cases Long et al.
+(2023); Tu et al.
+(2023); Joshi et al.
+(2024).
+Long et al.
+(2023) propose strategies for amending LLMs’ output based on consistency properties in causal inference.
+Feng et al.
+(2024) found the LLMs’ performance on causal discovery tasks depends on pretraining corpora (corpora with higher frequency of causal mentions perform better) and provided context when making the predictions.
+To alleviate the impact of erroneous causal information from LLMs, previous works have integrated LLMs with traditional causal discovery methods.
+Causal discovery or causal structure learning is the task of recovering causal graphs from observational data whenever possible Zanga et al.
+(2022).Vashishtha et al.
+(2023) propose two algorithms that combine LLMs with causal discovery methods: the first uses causal order from an LLM to orient the undirected edges outputted by a constraint-based algorithm and the second utilizes the LLM causal order as a prior for a score-based algorithm.
+Similarly, Li et al.
+(2024) introduced LLM-guided meta-initialization to extract the meta-knowledge from textual information to improve the quality of a temporal causal discovery method.
+Khatibi et al.
+(2024) used LLMs to refine the output of data-driven causal discovery algorithms.
+Ban et al.
+(2023b) incorporated LLM-driven causal statements as qualitative ancestral constraints in the Bayesian network structure to guide data-driven algorithms.
+Ban et al.
+(2023a) then propose an iterative framework to validate and finetune based on LLM feedback.
+For a broader discussion on integrating LLMs into causal discovery methods, we refer readers to a recent survey on this specific topic Wan et al.
+(2024).
+4.3 Future Directions
+LLMs can significantly contribute to overcoming the current limitations of causal inference methods as a general expert.
+A common assumption in many causal methods is the existence of corresponding data points for every treatment.
+However, this assumption often proves untrue, particularly when dealing with imbalanced minority data that may not support meaningful learning.
+LLMs, functioning as versatile experts, have the potential to address this challenge by aiding in data augmentation for minority data.
+Through their comprehensive understanding of language and context, LLMs can enhance the availability of diverse data points, facilitating more robust and effective causal inference in situations where traditional methods may struggle due to data imbalances.
+Similarly, many methods operate under a strong assumption of unconfoundedness within the potential outcome framework Historically, this assumption has been accepted due to a lack of domain knowledge regarding the underlying causal graph or identification of potential confounders.
+However, LLMs offer an opportunity to alleviate the limitation.
+5 Conclusion
+At its core, a large language model (LLM) is like a vast library of knowledge.
+One of the ongoing challenges is figuring out how to extract and use this knowledge effectively.
+The key to improving LLMs lies in enhancing their ability to understand cause and effect – essentially, how things are connected.
+Causal reasoning is crucial for making LLMs smarter.
+Looking at it from a causal inference perspective, we find a valuable framework that helps boost the effectiveness of LLMs.
+Meanwhile, as keepers of human knowledge, LLMs can even help overcome limitations in causal inference by providing broad expertise that goes beyond existing constraints, reshaping our understanding in this important area and bringing new vitality to this area.
+In this survey, we offer a thorough examination of both directions and concise summaries of the methods scrutinized, offering a comprehensive overview of the current state of research at this intersection.
+Limitations
+With the increasing popularity of large language models (LLMs), understanding their reasoning abilities becomes ever more crucial.
+Many tasks performed by LLMs require an understanding of causality, making the evaluation and enhancement of their reasoning and causal inference abilities a key focus.
+This is where causal inference plays a critical role.
+At the same time, causal inference itself relies on a certain level of world knowledge, which LLMs are well-suited to provide.
+In this survey paper, we aim to provide a comprehensive review of how LLMs contribute to causal inference and, in turn, how causal inference can improve LLMs.
+However, the intersection of causal inference and LLMs represents a rapidly evolving research frontier where significant developments appear across NLP/ML venues and preprints.
+Given this field’s dynamic nature, our selection methodology prioritizes comprehensive coverage of key developments over exhaustive enumeration.
+For preprints, the authors of this paper manually reviewed them to assess their quality and relevance to the topic.
+There are several possible ways to organize this paper, particularly for Section 3.
+We follow the framework proposed by Feder et al.
+(2022) on how causal inference can enhance NLP models, focusing on performance, robustness, fairness, and interpretability.
+While Feder et al.
+(2022) focuses on embedding-based methods, our survey reflects the transformative impact of LLMs with strong generative capabilities.
+We supplement the categorization with additional discussions on LLM’s reasoning capacity and multi-modality, capturing research trends driven by the enhanced capacities of LLMs.
+It’s worth noting that some papers may fit into multiple sections, and we grouped them based on our judgment of the papers’ main contributions.
+Appendix A Brief Introduction of Causal Inference
 In this section, we present the background knowledge of causal inference, including task descriptions, basic concepts and notations, and general solutions.
@@
 Therefore, estimating the treatment effect from observational data has attracted growing attention due to the wide availability of observational data, and methods are developed for the investigation of the causal effect of a certain treatment without performing randomized experiments.
-3.1 Potential Outcome Framework
-One of the most influential frameworks in identifying and quantifying causal effects in observational data is the potential outcomes framework [80].
+A.1 Potential Outcome Framework
+One of the most influential frameworks in identifying and quantifying causal effects in observational data is the potential outcomes framework Rubin (1974).
 The potential outcomes approach associates causality with manipulation applied to units, and compares causal effects of different treatments via their corresponding potential outcomes.
-Following [80], we state basic concepts in the potential outcome framework.
+Following Rubin (1974), we state basic concepts in the potential outcome framework.
 Unit.
 A unit is the atomic research object in the treatment effect study.
-A unit can be a physical object, a firm, a patient, a person, or a collection of objects or persons, such as a classroom or a market, at a particular time point [80].
+A unit can be a physical object, a firm, a patient, a person, or a collection of objects or persons, such as a classroom or a market, at a particular time point Rubin (1974).
 Under the potential outcome framework, the atomic research objects at different time points are different units.
@@
 At the subgroup level, the treatment effect is called the Conditional Average Treatment Effect (CATE).
-Definition 3.1 (Binary Average Treatment Effect(ATE)).
+Definition A.1 (Binary Average Treatment Effect(ATE)).
 Suppose we want to measure the treatment effect of a treatment $T=1$. Then the average treatment effect is defined as:
@@
 where $Y(T=1)$ and $Y(T=0)$ denote the potential treated and control outcome of the whole population respectively.
-Definition 3.2 (Conditional Average Treatment Effect (CATE)).
+Definition A.2 (Conditional Average Treatment Effect (CATE)).
 $\mathbb{E}[Y(T=1)|X=x]-\mathbb{E}[Y(T=0)|X=x]$ (5)
 where $\mathbb{E}[Y(T=1)|X=x]$, $\mathbb{E}[Y(T=0)|X=x]$ are the potential treated and control outcome of the subgroup with $X=x$.
-At the individual level, the treatment effect is defined as Individual Treatment Effect (ITE). In some literature, ITE is treated as the same as CATE [75].
-3.2 Causal Graphical Models
+At the individual level, the treatment effect is defined as Individual Treatment Effect (ITE). In some literature, ITE is treated as the same as CATE Pearl (2009).
+A.2 Causal Graphical Models
 The potential outcome framework is powerful in recovering the effect of causes.
@@
 In the front of the challenge, causal graphical models utilize directed edges to represent causalities and encode conditional independence among variables in the graphs.
-3.2.1 Structural Equation Models (SEMs)
-One of the most widely-spread formulations is the Structural Equation Model [99, 76], where linear structural equation models are used to present causal relationships by directed edges, which differentiate correlation from causation when the graph structure is given.
-The linearity assumption was later been relaxed by [76] and it formalized causal graphical models for presenting causal relations using Directed Acyclic Graphs (DAGs).
+A.2.1 Structural Equation Models (SEMs)
+One of the most widely-spread formulations is the Structural Equation Model Wright (1934); Pearl (1998), where linear structural equation models are used to present causal relationships by directed edges, which differentiate correlation from causation when the graph structure is given.
+The linearity assumption was later been relaxed by Pearl (1998) and it formalized causal graphical models for presenting causal relations using Directed Acyclic Graphs (DAGs).
 Specifically, consider the random variable $\mathbf{X}\in\mathcal{R}^{D\times N}=[X_{1},X_{2},...,X_{N}]$, the linear SEM consists of a set of equations of the form:
@@
 The random variables $\mathbf{X}$ that satisfies the model structure of the form in Equation 6 or Equation 7 can be represented by a directed acyclic graph (DAG) $G=(V,E)$, where $V$ is the set of associated vertices, each corresponding to one of a variable of interest $X_{i}$, and $E$ is the corresponding edge set.
-With pre-specified DAG and assumptions on the latent variables, the coefficients between the latent variables are identifiable[46].
-3.2.2 Bayesian Network
+With pre-specified DAG and assumptions on the latent variables, the coefficients between the latent variables are identifiableKuroki and Pearl (2014).
+A.2.2 Bayesian Network
 Causal inference can be naturally embedded in graphical model frameworks since the dependencies and interactions between variables can be presented by graphs with probabilistic distributions, in which nodes correspond to variables of interest and edges represent associations.
@@
 In the next section, we show a comprehensive survey of how existing works help with the tasks and challenges in LLMs in detail.
-4 Causal Inference for Large Language Models
-LLMs can significantly benefit from causal inference as it enhances their ability to understand and reason about cause-and-effect relationships within data.
-In this section, we review how LLMs can benefit from a causal view in the following perspectives, understanding and improving the LLMs’ reasoning capacity (Section 4.1), addressing fairness (Section 4.2) and safety (Section 4.3) issues in LLMs, complementing LLMs with explanations (Section 4.4), and handling multimodality (Section 4.5).
-We then organize benchmark datasets from these perspectives in Section 4.6.
-4.1 Reasoning Capacity
-4.1.1 Model Understanding
-LLMs have demonstrated many emerging abilities in language generation and certain reasoning tasks [14, 44].
-As the reasoning process is often associated with causal factors, it is logical to first understand and evaluate LLMs’ reasoning ability from a causal lens.
-Zečević et al.
-[109] argued LLMs are not causal and hypothesized that LLMs are simply trained on the data, in which causal knowledge is embedded.
-Thus, in the inference stage, the LLMs can directly recite the causal knowledge without understanding the true causality in the context.
-Similar behaviors are exhibited in a Causal Reasoning Assessment Benchmark, CRAB, that consists of 1.8K causal frames and 352 causal chains in real-world narratives [79], where LLMs are required to output the causality class (high, medium, low, or no causality) between variables.
-They show that LLMs can capture explicit causal statements in pre-training data, but face performance drop when applying causal reasoning to new distributions, i.e., events that happened after the pre-training phase.
-Kim et al.
-[43] examined LLMs’ abilities in understanding the causalities of both scientific papers and newspapers.
-Their evaluation protocol is designed for the LLM to tell whether a statement is causal, conditional causal, correlational, or no relationship.
-The results show that ChatGPT performs worse than a fine-tuned BERT model in understanding causality.
-Abdali et al.
-[1] show the effectiveness of applying LLMs to diagnose the cause of issues from Microsoft Windows Feedback Hub.
-Li et al.
-[47] showed that LLMs can identify dynamical (spatio-temporal) effects.
-However, how to infer the relationship and interactions of them is still challenging for LLMs, which are more emphasized as causal structures in causal inference.
-Another important line of work is to understand LLMs’ hallucination and faithfulness in knowledge reasoning by considering causal effects.
-Tang et al.
-[88] proposed a multi-agent system, CaCo-CoT, where some LLMs are reasoners and others are evaluators.
-Reasoners try to provide causal solutions, while evaluators try to challenge the reasoners with counterfactual candidates.
-With the cooperative reasoning framework, CaCo-CoT helps to improve causal-consistency.
-Zhang et al.
-[116] identified the potential knowledge bias pretrained in the LLMs as the confounder which causes incorrect answers and hallucinations.
-Zhang et al.
-[116] proposed a chain-of-question framework to generate sub-questions necessary to answer a question, and involve humans in the loop to provide the correct answers.
-With human annotation in the loop, the confounding causal effect can be reduced, thus mitigating the spurious correlation.
-4.1.2 Commonsense Reasoning
-Commonsense reasoning involves the ability to apply everyday knowledge and intuitive understandings of the world to make decisions or draw conclusions, which is vital for LLMs’ contextual understanding and human-like interactions [21, 84].
-This section briefly summarizes the commonsense reasoning ability of LLMs under various settings [27, 98] and the employment of causally motivated methods in improving commonsense causality reasoning [114, 123, 62, 17, 120].
-Through the evaluation of ChatGPT’s performance on event causality identification, causal discovery, and causal explanation generation, Gao et al.
-[27] have shown that ChatGPT is not a good reasoner but a good causal explainer.
-ChatGPT, and even gpt-4 is outperformed by baseline methods based on fine-tuned small pre-trained language models on event causality identification and does poorly on causal discovery.
-They also observe serious hallucinations on causal reasoning under In-Context learning and Chain-of-Thought settings A similar conclusion has also been obtained in [98], which analyzed the causal question answering capabilities of LLMs.
-It concluded that LLMs did not arrive at their answers through reasoning but through memorization of the corresponding question and answer pair.
-Even if ChatGPT may not have causal reasoning ability, it generates accurate and detailed causal explanations in some cases [27].
-To improve the commonsense causality reasoning of LLMs that identify causes from effects in natural language, ROCK (Reason O(A)bout Commonsense K(C)causality) [114] balances confounding effects using temporal propensities.
-The central question in this framework is the estimation of Average treatment effect (ATE).
-Given two events such that $E_{1}$ precedes $E_{2}$, the strength of causation from $E_{1}$ to $E_{2}$ can be estimated by the change in the likelihood of $E_{2}$’s occurrence due to intervening on $E_{1}$, denoted as ATE.
-While ROCK adopts a potential-outcome framework, Chen et al.
-[17] uses a conversation cognitive model based on intuition theories, and transforms intuitive reasoning into a structural causal model.
-Aiming to improve conversation reasoning, the authors incorporate mental states such as desires, memory, experience and emotion, as an unobservable exogenous variable that implicitly affects the corresponding observable utterance:
-$H=AH+E,$ (6)
-where $H$ is the embedding of the utterance, $E$ is the exogenous variable, and $A$ is the adjacency matrix.
-Other than facilitating reasoning ability of LLMs directly, Zheng et al.
-[123] use causal inference to preserve commonsense knowledge from Pre-trained language models for fine-tuning to prevent catastrophic forgetting; Lu et al.
-[62] focus on improving LLM’s ability in generalized procedural planning with commonsense-infused prompts.
-In particular, Zheng et al.
-[123] abstract the fine-tuning process as a causal graph and discovered that catastrophic forgetting is due to missing causal effects from the pre-trained data.
-To preserve old knowledge from a pre-trained language model, an objective for fine-tuning (causal effect tuning) is introduced.
-For procedural planning tasks, Lu et al.
-[62] proposed to learn causeeffect relations among complex goals and stepwise tasks, and reduced spurious correlation among them via front door adjustment.
-4.1.3 Counterfactual Reasoning
-Another potential use of LLMs’ generative capacity is to generate counterfactuals for data-augmentations for small language models.
-Given a text $x$ and a black-box classifier $B$, the counterfactual text $\tilde{x}$ of text $x$ should satisfy the following requirement [11, 70]:
-  1. 1.
-    $\tilde{x}$ has a different class than $x$, $B(x)\neq B(\tilde{x})$.
-  2. 2.
-    $x$ and $\tilde{x}$ differ only by minimal lexical changes.
-  3. 3.
-    $\tilde{x}$ is a feasible text and the commonsense constraint is satisfied.
-This section discusses briefly the performance of LLMs in generating counterfactuals [50, 60] and the effort of improving their qualities [70].
-Li et al.
-in [50] examined the effectiveness of LLMs’ performance in generating counterfactuals under four tasks: (1) sentiment analysis (SA) that alters sentiment polarity; (2) natural language inference (NLI) that given a premise sentence and a hypothesis sentence, make a change to the hypothesis sentence to alter the relationship between it and the premise sentence; (3) named entity recognition (NER) that changes the entities in a sentence whose type is the same as the original entity type; (4) relation extraction (RE) that changes the relation between the head and tail entity.
-It has been demonstrated that for simple tasks like SA and NLI, data augmented via LLMs can mitigate potential spurious associations.
-For more complicated tasks like RE, LLMs may generate low-quality counterfactuals.
-In order to generate high-quality counterfactual, detailed instructions are crucial.
-However, counter-intuitively, chain-of-thought doesn’t always help as it may even lead to significant performance decreases under some settings.
-On the other hand, Liu et al.
-in [60] evaluated abductive reasoning and counterfactual reasoning abilities for large language models of code (Code-LLMs) and compared them with text models.
-With code-prompts designed to tackle causal reasoning tasks, it has been shown in [60] that Code-LLMs achieve better results than text models.
-As shown in the evaluation [50], though LLMs as counterfactual generators can enhance the performance of small language models on simple tasks such as sentiment analysis and natural language inference, the generated counterfactuals fail to have any significant effect on complex tasks like relation extraction.
-Miao et al.
-in [70] claim that this is due to the failure of identifying causal terms correctly and ignoring the commonsense constraint.
-To amend this, they proposed a framework to generate commonsense counterfactuals for stable relation extraction via an intervention-based strategy.
-This framework is demonstrated to have enhanced the stability on relation extraction tasks under various settings including the low-resource, out-of-domain and adversarial-attack scenarios.
-4.2 Fairness and Bias
-Fairness and bias are pivotal factors in deploying language models effectively and ethically.
-Bias is common in pretrained language models as they capture and potentially amplify undesired social stereotypes and biases [68, 2, 126, 100, 96].
-An example of bias in language models includes gender associations with specific professions, such as male firefighters and female nurses [82].
-Causality-based methodologies offer a promising approach for mitigating biases in language models by discerning the origins of bias through a causal perspective.
-Bias mitigation is then followed by eliminating the unwanted spurious correlation between generative factors though different types of causal intervention or causal invariant learning.
-Ding et al.
-[22] introduce a proxy variable related to gender bias in the causal graph, and use two different ways to eliminate the potential proxy bias and unresolved bias under the linear structural equation model.
-Zhou et al.
-[124] believe that the backdoor path between the ground truth label and the non-causal factors is the source of bias, and uses the Independent Causal Mechanism (ICM) principle to mitigate bias.
-Their proposed method, Causal-Debias, achieves causal intervention and creates interventional distributions with respect to different demographic groups by augmenting and expanding the original data distribution.
-Wang et al.
-[95] raise concerns over the precision of parameter estimation in existing causal models and introduce several intermediate variables that are the causal children of raw text and the causal parent of the input to language models.
-Under this assumption, they propose to eliminate bias by performing a ‘do’ operation on the intermediate variables for both white-box LLMs and black-box LLMs.
-Madhavan et al.
-[65] consider the tokens generated by generative language models trained with causal language modeling (CLM) objectives as a causal graph, and analyze the bias under this model.
-Jenny et al.
-[37] use Activity Dependency Networks (ADNs) to describe the causality effect between normative variables, such as clarity and authenticity, to structure the cause of bias.
-The authors argue that using ADNs can better explain previously simplified views of bias using just correlation, and display the complexity nature in identifying and mitigating biases in large language models.
-4.3 Safety
-With the application of language models in multiple downstream tasks, researchers have observed the unreliability phenomenon of LLMs in knowledge probing [15] or downstream inference tasks [9, 122, 118, 69].
-There is increasing interest in applying the causal inference technique to analyze the causality of the non-robustness of the model and adjust the treatment to resolve the challenges [15, 118, 122, 56].
-Knowledge probing
-To elicit the knowledge encoded in LLMs, previous work uses prompt-based prompting, that is, querying LMs with task-specific prompts.
-However, in this process, LLMs face challenges of unreliability [15, 102], such as using shortcuts to complete the probing and generating different predictions for the semantically equivalent prompt.
-To explore the reason behind the non-robustness, an empirical study from [15] constructs a structural causal model (SCM) containing 11 variables and identifies three types of bias in the knowledge probing procedure: prompt preference bias, instance verbalization bias, and sample disparity bias.
-Blocking the corresponding backdoor path for each type of bias effectively eliminates the bias.
-This finding reveals the potential of constructing an SCM with LLM-related variables to improve LLM robustness.
-Downstream Inference Tasks
-In addition to knowledge probing, LLMs show vulnerability when encountering attacks on the prompts in downstream inference tasks [118, 122].
-By simply translating text tokens in input prompts to emoji sequences, LLMs generate more severe hallucinations [122].
-Some neurons within LLMs also have an unreasonably high causal effect on the model response and, by changing the value of that neuron, LLMs will produce meaningless responses [122].
-The main reason why LLMs fail in prompt attacks is that it uses the spurious correlation to make an inference [118].
-Training LLMs to learn the causal relationship between input $x$ and output $y$ is an intuitive method of better resisting prompt attacks.
-The randomized smoothing technique [112, 39] can model the interventional distribution $p(y|do(x))$ by assuming discrete adversarial perturbations as the Gaussian distribution [118].
-The method of smoothing in the latent semantic space is more robust against known attacks such as word substitutions, paraphrasing, and token position change [118].
-Causal inference techniques are useful to induce robustness for LLM applications.
-For the setting without prompt attacks, mitigating LLM unreliability can be achieved by identifying the causality and blocking the backdoor path.
-Moreover, smoothing the latent semantic space is effective in resisting prompt attacks.
-Despite the progress of causality analysis, there are some directions to explore, such as the causal relationship between the pretraining corpus and model robustness and the evaluation method of robustness in the generation task.
-4.4 Explainability
-Explainability in LLMs refers to the capacity to elucidate how these models arrive at their conclusions, enhancing transparency and trustworthiness in AI decision-making processes [32, 54].
-Many work have tried to explain and understand the inner workings of LLMs [35, 10, 33, 119, 55].
-We summarize research efforts that probe the causal mechanism in LLMs from the following three directions: intervening the inputs or prompts, intervening inner components of LLMs, and abstracting the working mechanism into a causal graph.
-Inputs or Prompt Intervention
-Input intervention, as a data-centric method, is to create counterfactual input text by changing the treated feature in the text and to observe the model behavior on the original and counterfactual texts.
-Recruiting individuals to produce counterfactual texts typically incurs significant expenses.
-However, the advent of LLMs has shown that generating counterfactual inputs can be achieved with less cost.
-LLMs can first identify the features in input texts causally associated with the predictions and are capable of changing the identified features to create the counterfactual texts [12, 28].
-These counterfactual texts can be utilized to investigate the causality of the LLM and can serve as a training dataset to learn a matching model, where the matched counterfactual pairs have similar embeddings [28].
-Various works have developed different prompting methods and found whether the prompting methods are causally associated with the final output of LLMs [81, 125, 97].
-However, the causal effect of prompting methods, such as chain-of-thought (CoT), and the final output is ambiguous.
-Prompt intervention, which alters only one particular aspect of prompts, is proposed to understand the contributions of each component of prompts on model behavior [64].
-The experiments from [64, 38] first find that the linguistic features and the grammar have a large effect on the LLM outputs.
-Then, intervention in intermediate variables in prompts leads to consistent final answers with the expected output of the hypothesized causal model [87].
-These findings suggest that LLMs realize the causal model suggested by their CoTs to a high extent, but LLMs also utilize spurious correlations such as sentence length to make responses.
-Input or prompt interventions are data-centric methods to probe the LLM mechanism, applied in both open-source or black-box LLMs.
-However, the detailed information cascade inside the LLMs cannot be discovered by such methods, so intervention on the inner model components is proposed.
-Inner Component Intervention
-The attention mechanism and multilayer perceptron (MLP) layers are the essential components in the structure of state-of-the-art (SoTA) LLMs.
-Existing work exchanges the activation values in MLP and attention layers of different inputs to probe the function of MLPs and attention mechanism in generating the answers [83].
-The experimental results indicate that LLMs transfer the information of inputs from midsequence early layers to the final token using the attention mechanism.
-Due to the complexity of LLMs, current work focuses only on math word problems with four fundamental arithmetic operators [83].
-It is an interesting direction to generalize the component intervention to other downstream tasks.
-Causal Graph Abstraction
-An intuitive way to characterize causality within LLMs is to abstract the working mechanism of LLMs into a causal graph.
-Boundless Distributed Alignment Search (DAS) [101] by replacing brute-force searching the original DAS [29] with learnable parameters, has been effective on the Alpaca model [89].
-Given four pre-defined causal models, Boundless DAS extracts two of them as the accurate hypotheses as the abstracted causal graph of the Alpaca model.
-However, the Boundless DAS method is restricted by the given causal hypothesis, and the future direction can explore how to abstract the causal graph in LLMs without prior causal graphs.
-Current explainability work from a causal view use LLMs to interpret the causality in real-world events.
-LLMs can generate high quality counterfactuals.
-By altering the inner values of LLMs and abstracting the causality in LLMs, the current work has pointed out a direction to characterize the inner causality of LLMs on various tasks.
-However, with these mentioned works in LLM interpretation, the scope of interpretation focuses on tasks that have a clear causal graph between task inputs and outputs such as simple math word problems.
-Probing the causality in LLMs for more complex generation tasks, such as question answering or summarization tasks, could provide more insight on the inner mechanism of LLMs.
-4.5 Multi-modality
-Large vision-language models have been popular in many applications [63, 48, 59, 30].
-How to conduct causal reasoning on both images and texts can be crucial in correctly answer multimodal questions.
-Pawlowski et al.
-[74] examined LLMs’ causal reasoning abilities and showed that the causal knowledge in the language models can be too strong a prior which often causes the model to neglect visual evidence.
-Ko et al.
-[45] proposed to alleviate the problem by adding self-consistent generation prediction, in which the three input V, Q and A are individually predicted based on the other two inputs.
-Specifically, Li et al.
-[49] proposed a image generation framework with causal reasoning and created a novel VQA datasets whose questions requires causal explanations.
-Another important question is to understand the spatial-temporal causal relationship of the visual elements within the images and videos.
-Su et al.
-[85] proposed a framework, CaKE-LM, to leverage the pretrained causal knowledge in the language model to understand the causal relations of the events in a video.
-Based on the generated causal reasoning results, CaKE-LM can further generate the question-answer pairs and construct a new benchmark for causal video question-answering.
-Tai et al.
-[86] proposed the link-context learning method to strengthen the LLM’s in-context learning abilities by instructing the model to understand the underlying causal relationship between the demonstration data points.
-Zhao et al.
-[121] suggested that there are two types of causal relationships in VQA.
-They proposed a prompting method, causal context generation, to engage contextual information for better VQA precision.
-4.6 Evaluation and Benchmark
-In this section, we list existing evaluation metrics and benchmarks from a causal perspective for LLMs as listed in Table 1.
-The causal evaluation mainly focuses on three aspects: the Model Understanding (MU) ability, the Commonsense Reasoning (CR) ability, the Counterfactual Reasoning (CF) ability, and the Fairness/Debias (FD) ability.
-Reference MU CR CF FD Language Multimodal
-ECHo [103]
-CREPE [115]
-CLOMO [36]
-IfQA [107]
-Cladder [41]
-MoCa [72]
-CORR2CAUSE [40]
-CVidQA [85]
-VQAI [49]
-Chen et al. [18]
-Gao et al. [27]
-CRAB [79]
-HELM [53]
-Fair-Prism [25]
-Biasasker [94]
-Table 1: We summarize the existing evaluation benchmarks.
-Based on the evaluation tasks, we categorize the benchmarks into three categories: Model Understanding (MU), Commonsense Reasoning (CR), Counterfactual Reasoning (CF) and Fairness/Debias (FD).
-Based on the modalities of the data samples, we identify the benchmarks with only textual inputs (Language) and those with multimodal inputs (Multimodal).
-The benchmarks in model understanding (MU) focus on evaluating and understanding existing LLMs’ causal reasoning abilities in both natural language [41, 72, 40, 27, 79] and vison-language [85, 49, 103].
-In addition, some benchmarks [72, 27] also provide model understanding in comparison with human causal reasoning and moral judgments.
-Commonsense reasoning benchmarks (CR) evaluate LLMs on tasks that require extensive commonsense knowledge for both textual-only context [115, 41] and multimodal context [103].
-Contexts with commonsensical and anti-commonsensical are constructed in [41], to further investigate whether LLMs use averaged-out causal reasoning.
-Evaluating LLMs’ counterfactual reasoning (CF) abilities is essential in enabling explainable model reasoning and calibration of the generated rationales.
-Huang et al.
-[36] introduce a specific task and benchmark for assessing LLMs’ logical counterfactual thinking.
-Yu et al.
-[107] contribute a novel dataset to challenge LLMs in counterfactual reasoning in an open-domain QA context.
-Chen et al.
-[18] investigate the ability of LLMs to provide explanations that aid in understanding their reasoning process, particularly in the context of counterfactual scenarios.
-The fairness and bias (FD) evaluations are particularly in addressing biases, fairness, and the overall transparency of language models.
-HELM [53] is a comprehensive evaluation benchmark including previously neglected areas for fairness.
-Fair-Prism [25] focuses specifically on fairness-related harms in models, which are identified and measured by detailed human annotations.
-Biasasker [94] presents an automated framework to identify and measure social biases by probing the models with specially designed questions.
-5 Large Language Model for Causal Inference
-Causal inference, serving as a potent tool for addressing challenges in LLMs, heavily relies on world knowledge.
-As previously mentioned, there are three primary origins of causal inference: the potential outcome framework, graph-based causal methods, and the structural equations community.
-The potential outcome framework relies significantly on several assumptions to facilitate the comparison of treatment effects among groups/individuals.
-One of the most challenging aspects of applying the potential outcome framework lies in ensuring that these assumptions hold true in reality.
-In this section, we first examine these assumptions and subsequently illustrate how they are relaxed in existing literature.
-The graph-based causal methods and structural equation models also necessitate a certain level of understanding of the underlying causal graph.
-For instance, Directed Acyclic Graphs (DAGs) serve as a fundamental assumption, and many structural equation models assume a degree of linearity or that the input distribution adheres to specific probability distributions.
-In our review, we also explore how existing methods verify the distribution in the input data and extend current methodologies to accommodate more complex distributions with the assistance of LLMs.
-5.1 Fundamental Assumptions in Estimating Treatment Effect
-In existing causal inference literature, several assumptions are adopted to estimate the treatment effect.
-Here we discuss the three most commonly used assumptions and then show how the development of large language models could help relax or challenge these fundamental assumptions.
-Assumption 5.1 (Stable Unit Treatment Value Assumption ).
-The potential outcomes for any unit do not vary with the treatment assigned to other units, and, for each unit, there are no different forms or versions of each treatment level, which leads to different potential outcomes.
-This assumption emphasizes the independence of each unit when estimating the treatment effect that the units do not interact with each other.
-From a statistical perspective, it is equivalent to assuming each treatment assignment subject is iid.
-Assumption 5.2 (Ignorability/Unconfoundedness ).
-Given the observable background variable, X, treatment assignment T is independent of the potential outcomes.
-The 5.2 states that if the background variable $X$ is the same for two patients, then (1) the treatment assignment should be the same.
-(2) the potential outcome should also be the same.
-In other words, these two patients are treated as identical units, thus can be used to estimate the treatment effect if they are assigned different treatments in the static dataset, as the treatment assignment is treated as random.
-Assumption 5.3 (Positivity ).
-For any value of X, treatment assignment is not deterministic. i.e.,
-$P(T=t|X=x)>0,\forall t,x$ (7)
-This assumption tries to guarantee that the treatment effect can be estimated and that we can always find a comparable sample.
-In the binary case, if the task is to estimate the performance of a specific treatment $T=1$, we would need to compare the potential outcome of patients receiving the treatment against those who are not treated, requiring dataset points in both cases.
-5.2 Treatment Effect Estimation
-One main obstacle of traditional causal methods is the lack of counterfactual data, making the estimation of causal effects a difficult problem in practice.
-Chen et al.
-[19] proposed a new method for automatically generating high-quality counterfactual data at scale called DISCO (DIStilled COunterfactual Data).
-Specifically, it prompts to generate of phrasal perturbations with a large general language model.
-Then, a task-specific teacher model filters these generations to distill high-quality counterfactual data.
-In addition, Feder et al.
-[24] apply treatment effect estimation to align knowledge for generalization towards different domains.
-Zhang et al.
-[113] try to optimize the treatment effect estimation on unlabeled datasets by performing self-supervised causal learning through LLMs.
-Through exploring the primal-dual connection between optimal covariate balancing and self-attention, their method facilitates zero-shot causal inference through the final layer of a trained transformer-type architecture, contributing to a foundation model for treatment effect estimation.
-5.3 Causal Relationships Discovery
-Discovering causal relationships between variables is a fundamental step in causal inference as it enables the identification and estimation of causal effects.
-In this section, we introduce papers discussing how LLMs can help discover causal relationships.
-One line of this work focuses on casual relationship extraction or causality extraction which extracts causal relationships between two variables from text directly.
-Traditional methods rely on linguistic cues such as causal connectives (e.g., “cause”, “because”, and “lead to”) and grammatical patterns to identify causal pairs [104].
-A later work utilizes the power of statistical machine learning and deep learning to tackle this task in a supervised learning setting [105].
-As LLMs show promising potential with reasoning capacities as introduced in Section 4.1, many works use LLMs as a query tool to determine the edge direction between two given variables.
-For example, Kıcıman et al.
-[44] show that LLMs can achieve competitive performance in determining such pairwise causal relationships with accuracies up to $97\%$.
-Analyses in the medical domain [71, 4, 5] exhibit similar observations.
-However, other studies highlight LLMs’ limitations of such pairwise causal relationships.
-For example, sensitivity to prompt design leads to inconsistent results [61]; pairwise judgments can lead to cycles in the full causal graph [92]; pairwise judgments require large computational cost when applying to a large-scale dataset, $N$ variables would require $\binom{N}{2}$ prompts [7]; LLMs still provide false information despite achieving strong results in most cases [61, 91].
-Long et al.
-[61] propose strategies for amending LLMs’ output based on consistency properties in causal inference [61].
-To alleviate the impact of erroneous causal information from LLMs, previous works have integrated LLMs with traditional causal discovery methods.
-Causal discovery or causal structure learning is the task of recovering causal graphs from observational data whenever possible [108].
-Traditional causal discovery methods mainly include constrained-based methods that exploit a sequence of statistical tests and score-based methods that structure around the maximization of the fitness of a graph through a space of possible graphs.
-Vashishtha et al.
-propose two algorithms that combine LLMs with causal discovery methods: the first uses causal order from an LLM to orient the undirected edges outputted by a constraint-based algorithm and the second utilizes the LLM causal order as a prior for a score-based algorithm [92].
-Ban et al.
-incorporate LLM-driven causal statements as qualitative ancestral constraints in the Bayesian network structure to guide data-driven algorithms [8], which benefits smaller-scale problems, but encounters difficulties with larger datasets due to inaccuracies in the LLM-derived constraints.
-They then propose an iterative framework that employs LLMs to validate the accuracy of edges in the learned causal graph and fine-tune the causal discovery process based on LLM feedback [7].
-6 Future Directions
-Theoretical understanding of LLM’s reasoning capacity.
-Causal inference methods offer a promising avenue for gaining deeper insights into the reasoning capacity of Large Language Models (LLMs).
-A potential approach involves employing treatment effect estimation techniques to assess their performance on specific tasks.
-The gold standard for treatment effect estimation is comparing potential outcomes across various treatments in a controlled experiment.
-However, the practical challenges of conducting such experiments in real-world scenarios make the interactive nature of LLMs an ideal candidate for investigation.
-Leveraging their inherent interactivity, researchers could explore quasi-experimental settings, utilizing the natural variations in responses to infer causal relationships.
-It is crucial to acknowledge the complexities associated with interpreting LLMs, potential biases in training data, and the intricate nature of language understanding tasks.
-Additionally, considering ethical implications and addressing biases in both training and evaluation data is essential when delving into the reasoning capacity of LLMs using causal inference methods.
-Efficient training and inference in LLMs.
-As model scales and training data continue to expand, the process of knowledge updating and inference in Large Language Models (LLMs) becomes increasingly resource-intensive.
-Consequently, it becomes imperative for us to devise methods that can efficiently and judiciously update the knowledge base of pre-trained models.
-In this context, causal inference methods can play a crucial role by offering guidance in quantifying efficiency.
-By establishing causal relationships between methods of interest and the existing knowledge collection, these methods can help assess the impact of different updating strategies.
-This approach not only addresses the resource challenges associated with knowledge updating but also contributes to a more nuanced understanding of the evolving knowledge landscape in LLMs.
-LLM-based counterfactual estimation and augmentation.
-As a general expert, Large Language Models (LLMs) can significantly contribute to overcoming the current limitations of causal inference methods.
-A common assumption in many causal methods, as indicated in 5.3, is the existence of corresponding data points for every treatment.
-However, this assumption often proves untrue, particularly when dealing with imbalanced minority data that may not support meaningful learning.
-LLMs, functioning as versatile experts, have the potential to address this challenge by aiding in data augmentation for minority data.
-Through their comprehensive understanding of language and context, LLMs can enhance the availability of diverse data points, facilitating more robust and effective causal inference in situations where traditional methods may struggle due to data imbalances.
-Similarly, many methods operate under the assumption of unconfoundedness (5.2) within the potential outcome framework, a condition considered quite strong.
-Historically, this assumption has been accepted due to a lack of domain knowledge regarding the underlying causal graph or identification of potential confounders.
-However, with the advent of Large Language Models (LLMs), there is an opportunity to alleviate this stringent limitation.
-LLMs can act as general experts, offering valuable information about potential causal graphs and knowledge.
-This transformative capability of LLMs opens avenues to enhance our understanding of causal relationships, addressing a historical challenge where unconfoundedness assumptions were made due to limited domain knowledge.
-7 Conclusion
-At its core, a large language model (LLM) is like a vast library of knowledge.
-One of the ongoing challenges is figuring out how to extract and use this knowledge effectively.
-The key to improving LLMs lies in enhancing their ability to understand cause and effect – essentially, how things are connected.
-Causal reasoning is crucial for making LLMs smarter.
-Looking at it from a causal inference perspective, we find a valuable framework that helps boost the effectiveness of LLMs.
-Meanwhile, as keepers of human knowledge, LLMs can even help overcome limitations in causal inference by providing broad expertise that goes beyond existing constraints, reshaping our understanding in this important area and bringing new vitality to this area.
-In this survey, we offer a thorough examination of the current landscape where large language models (LLM) intersect with causal inference.
-We delve into how causal inference contributes to LLM, enhancing aspects such as reasoning, fairness, and safety, along with the explainability of LLM.
-Additionally, we explore how LLM, in turn, broadens the horizons of causal inference.
-Across these categories, our survey provides in-depth discussions, comparisons, and concise summaries of the methods scrutinized, offering a comprehensive overview of the current state of research at this intersection.
-The available benchmark datasets and open-source codes of those methods are also listed.
-The examination of the current advancements in causal inference and large language models serves a dual purpose.
-Firstly, it enhances our comprehension of how these two fields mutually benefit from each other.
-Secondly, it catalyzes the emergence of novel questions, propelling us closer to achieving Artificial General Intelligence.
-Moreover, this exploration has the potential to extend into diverse realms and find applications in real-world scenarios, showcasing the far-reaching implications of the synergy between causal inference and LLM.